### v1

# Dynamic asymptotic dimension growth for group actions and groupoids

Abstract.
We introduce the notion of dynamic asymptotic dimension growth for actions of discrete groups on compact spaces, and more generally for locally compact étale groupoids with compact unit space. Using the work of Bartels, Lück, and Reich to bridge our notion for group actions with asymptotic dimension growth for groups, thus providing many concrete examples for our notion. Moreover, we show that the asymptotic dimension growth for a discrete metric space of bounded geometry is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid, hence deducing that the coarse groupoid with subexponential dynamic asymptotic dimension growth is amenable. More generally, every $\sigma$-compact locally compact Hausdorff étale groupoid with compact unit space having dynamic asymptotic dimension growth at most $x^{\alpha}$ $(0<\alpha<1)$ is amenable.
Key words and phrases:
Dynamic asymptotic dimension growth, Group actions, Groupoids, Amenability
2020 Mathematics Subject Classification:
54F45, 22A22, 37A55, 37B05, 46L05, 54H20
Hang Wang is supported by NSFC 12271165 and in part by the Science and Technology Commission of Shanghai Municipality (No. 22DZ2229014, 23JC1401900)
Yanru Wang is supported by Technology Commission of Shanghai Municipality (No.23JC1401900). Yanru Wang is the corresponding author.
Jianguo Zhang is partially supported by NSFC 12271165, 12171156, and 12301154
1. Introduction
The notion of asymptotic dimension was primarily proposed by Gromov in [Gro93] as a large-scale analogue of the Lebesgue covering dimension. The idea gained widespread attention following Yu’s work, demonstrating that the Novikov conjecture holds for groups with finite asymptotic dimension (see [Yu98]). This result motivated numerous mathematicians to explore groups and spaces with finite asymptotic dimension. For instance, Gromov initially established that hyperbolic groups and metric spaces have finite asymptotic dimension (see [Gro93]). Later, Roe provided a more straightforward proof of this finding (see [Roe05]). Following this, Bell and Dranishnikov showed that the free product, the amalgamated free product, and the HNN extension can preserve finite asymptotic dimension (see [BD01]). Four years later, Osin demonstrated that if a finitely generated group $G$ is hyperbolic relative to a collection of subgroups with finite asymptotic dimension, then $G$ itself has finite asymptotic dimension (see [Osi05]). Next, Wright proved that the asymptotic dimension of a finite-dimensional CAT(0) cube complex is bounded above by its dimension (see [Wri12]). Bestvina, Bromberg, and Fujiwara then proved that mapping class groups have finite asymptotic dimension (see [BBF15]). Following that, Behrstock, Hagen, and Sisto proved that all hierarchically hyperbolic groups have finite asymptotic dimension, thereby recovering results related to including mapping class groups, relatively hyperbolic groups, and CAT(0)-cube complexes (see [BHS17]), among many others. For a comprehensive survey of this theory, refer to [BD11]; for recent developments, see [FP21, Tse23, FS24]. Moreover, there are several interactions between this area and $L^{p}$ operator algebras. In the work of [ZZ21], Zhang and Zhou showed that the $L^{p}$ coarse Baum-Connes conjecture holds for proper metric spaces with finite asymptotic dimension when $p\in[1,\infty)$. More recently, Wang, Wang, Zhang, and Zhou demonstrated that any $L^{p}$ Roe algebra for a coarsely uniformly contractible discrete metric space of bounded geometry with finite asymptotic dimension satisfies the persistence approximation property for quantitative $K$-theory of filtered $L^{p}$ operator algebras (see [WWZZ24]).
However, there are still plenty of groups known to have infinite asymptotic dimension, such as the wreath product $\mathbb{Z}\wr\mathbb{Z}$ (see [Dra06]), the first Grigorchuk group (see [Smi07]), Thompson’s group (see [DS11]), Celebrated Gromov’s group in [Gro03] containing an expander. To study the growth of dimension functions of groups with infinite asymptotic dimension, Dranishnikov [Dra06] introduced the notion of asymptotic dimension growth of a metric space. This concept generalizes the notion of finite asymptotic dimension since if the asymptotic dimension function is constant, then the space is of finite asymptotic dimension. Dranishnikov proved that the wreath product of a finitely generated nilpotent group with a finitely generated group of finite asymptotic dimension exhibits polynomial asymptotic dimension growth. Furthermore, he showed that all finitely generated groups have asymptotic dimension growth at most exponential [Dra06]. Additionally, he proved that polynomial asymptotic dimension growth implies Yu’s property A, a weak amenability-type condition introduced by Yu [Yu00]. Combined with Yu’s theorem on the coarse Baum-Connes conjecture [Yu00] implies that spaces with polynomial asymptotic dimension growth satisfy the Baum-Connes conjecture, and hence the Novikov conjecture. Extending Dranishnikov’s result, Ozawa [Oza12] established that a metric space with subexponential asymptotic dimension growth has property A; For another reference, see [Opp14] for the case of discrete metric spaces of bounded geometry.
Many examples of groups and spaces are shown to have subexponential asymptotic dimension growth. An absorbing example is a geodesic coarse median space with finite rank and at most exponential volume growth [ANWZ18], which includes the hyperbolic space and the mapping class group. Another interesting example is the direct product defined as $G=\bigoplus\limits_{n\in\mathbb{N}}f(n)\mathbb{Z}$, where $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ is a strictly increasing function that grows slower than a linear function. When endowed with the direct sum metric, $G$ becomes a discrete metric space of bounded geometry with slow asymptotic dimension growth [CFY09], a specific case of subexponential dimension growth. Besides, Bell examined how various constructions involving groups affect the asymptotic dimension function [Bel05]. To apply this notion to coarse geometry, we will provide a couple of equivalent interlocking definitions of the asymptotic dimension growth in this paper. Moreover, the last equivalent definition lays a good foundation for the theorem in Section 4.
Dynamic asymptotic dimension, developed by Guentner, Willett, and Yu in their paper [GWY17], is a concept of dimension for actions of discrete groups on locally compact spaces, as well as locally compact Hausdorff étale groupoids. This notion is connected to the work of Bartels, Lück, and Reich [BLR08, Bar16] on the Farrel-Jones conjecture in the context of controlled topology and the nuclear dimension in the classification of $C^{*}$-algebras. Subsequently, they extended this idea to dynamical complexity, demonstrating that the Baum-Connes conjecture holds for actions exhibiting finite dynamical complexity [GWY22].
The main aim of this paper is to investigate the growth of dynamic asymptotic dimension function in the context of group actions and groupoids. This study is inspired by Guentner, Willett, and Yu’s work [GWY17], which explores finite dynamic asymptotic dimension in relation to dynamics, topology, coarse geometry, and $C^{*}$-algebras. Their research [GWY22] on dynamical complexity and controlled operator $K$-theory, as well as the work of [ANWZ18] Arzhantseva, Niblo, Wright, and Zhang on characterizing asymptotic dimension growth, forms a solid foundation for our study. These influential works provide valuable methods and results that facilitate our investigation.
In the case of group actions, we will define the dynamic asymptotic dimension growth for actions of discrete groups on compact spaces. This concept can be viewed as a generalization of the finite dynamic asymptotic dimension, and serves as the dynamical version of the asymptotic dimension growth. Analogous to the finite case, we can connect our notion to the work of Bartels et al. [BLR08, Bar16] on the Farrel-Jones conjecture to introduce a new notion called an $h$-BLR action, where $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ is a nondecreasing function. We will show that an $h$-BLR action has dynamic asymptotic dimension function at most $h$. Furthermore, when $h$ grows at a rate slower than $\sqrt{x}$, we derive an equivariant condition related to the above BLR condition, which is formally similar to the definition of asymptotic dimension growth.
In the case of groupoids, we will introduce the dynamic asymptotic dimension growth for locally compact Hausdorff étale groupoids with compact unit space. This notion generalizes the case of group actions. Moreover, we can relate our notion to coarse geometry. More precisely, for a discrete metric space $X$ of bounded geometry, we obtain the following theorem:
Theorem 1.1 (see Theorem 4.12).
Let $X$ be a discrete metric space of bounded geometry, and let $G(X)$ be the associated coarse groupoid. Let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a non-decreasing function. Then $\textnormal{ad}_{X}\approx f$ if and only if $\textnormal{dad}_{G(X)}\approx f$.
This theorem demonstrates that the asymptotic dimension growth for $X$ is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid $G(X)$. The critical point in our proof is to combine the last equivalent definition of asymptotic dimension growth with the proof method that appeared in [[GWY17], Theorem 6.4].
Combining this theorem with Ozawa’s result that the subexponential asymptotic dimension growth implies property A, we derive the following corollary:
Corollary 1.2 (see Corollary 4.15).
Let $X$ be a discrete metric space of bounded geometry. If the associated coarse groupoid $G(X)$ has subexponential dynamic asymptotic dimension growth, then $G(X)$ is amenable.
This corollary provides an alternative sufficient condition for the amenability of the coarse groupoid. For this result, we can give a concrete example: for a geodesic uniformly locally finite coarse median space with finite rank and at most exponential volume growth $X$, its corresponding coarse groupoid $G(X)$ has subexponential dynamic asymptotic dimension growth, and is thus amenable.
With the help of the above BLR condition, we can construct the equivalence between the asymptotic dimension growth for countable discrete groups and the dynamic asymptotic dimension growth for group actions. As a consequence, we have the following theorem:
Theorem 1.3 (see Theorem 4.18).
Let $\Gamma$ be a countable infinite discrete group with subexponential asymptotic dimension growth $h$. Then $\Gamma$ admits a free, minimal action on the Cantor set which has dynamic asymptotic dimension growth at most $h$.
This theorem is an adaption and a generalization of [[GWY17], Theorem 6.6]. Moreover, it provides a shortcut to construct more examples of group actions with subexponential dynamic asymptotic dimension growth.
The amenable groupoids behave well in several areas, such as the Baum-Connes conjecture [Tu99], the Novikov conjecture [Hig00], and the classification of $C^{*}$-algebras [AR00]. The main implication we explore in this paper is the amenability of groupoids. For this, we develop a key tool that is not only partitions of unity. Using this tool, we can directly prove the following main theorem:
Theorem 1.4.
(see Theorem 6.4) Let $\mathcal{G}$ be a $\sigma$-compact locally compact Hausdorff étale groupoid with compact unit space, and with dynamic asymptotic dimension growth $f\preceq x^{\alpha}(0<\alpha<1)$. Then $\mathcal{G}$ is amenable.
This theorem is a generalization of Corollary 1.2 in the case of slow growth. The idea is also inspired by Guentner, Willett, and Yu’s work [[GWY22], Appendix A ] on finite dynamical complexity for étale groupoids.
The outline of this paper is as follows. In Section 2, we give several equivalent definitions of the asymptotic dimension growth and provide several typical examples. In Section 3, we introduce the dynamic asymptotic dimension growth for group actions and discuss its connection to the work of Bartels, Lück, and Reich. In Section 4, we define the dynamic asymptotic dimension growth for groupoids. Furthermore, we show that the asymptotic dimension growth for a discrete metric space of bounded geometry is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid. This allows us to construct the equivalence between the asymptotic dimension growth for countable discrete groups and the dynamic asymptotic dimension growth for group actions. In Section 5, we develop a technical tool more than partitions of unity, which plays a vital role in proving amenability. Finally, in section 6, we find a sufficient condition for amenable groupoids: every $\sigma$-compact, locally compact Hausdorff étale groupoid $\mathcal{G}$ with compact unit space having dynamic asymptotic dimension growth at most $x^{\alpha}$ $(0<\alpha<1)$ is amenable.
2. Asymptotic dimension growth for metric spaces
The volume growth of finitely generated groups can be at most exponential. To explore the intermediate dimension growth of such groups, Dranishnikov in [Dra06] proposed the notion of the asymptotic dimension growth, which is a generalization of the finite asymptotic dimension. Additionally, he gave examples of finitely generated groups with polynomial asymptotic dimension growth and showed that all such groups have property A [Dra06]. Later, Ozawa extended this result to the metric spaces with subexponential asymptotic dimension growth [Oza12]. Combined with Yu’s result that the coarse Baum-Connes conjecture holds for spaces with property A, we see that this conjecture holds for spaces with subexponential asymptotic dimension growth. In this sense, subexponential asymptotic dimension growth is weaker than finite asymptotic dimension and stronger than property A. This section will give a couple of equivalent definitions of the asymptotic dimension growth. Moreover, we apply the last equivalent definition to coarse geometry and the proof of Theorem 4.12
2.1. Equivalent definitions
In this subsection, we will review the notion of the asymptotic dimension growth, and introduce two equivalent definitions. To see this concept more clearly, we will provide several concrete examples.
We review the following notions to define the asymptotic dimension function. Let $(X,d)$ be a metric space and $R>0$. Say that a family $\mathcal{W}$ of subsets of $X$ is $R$-disjoint if $d(W,W^{\prime})>R$ for every $W\neq W^{\prime}$ in $\mathcal{W}$. Say that the family $\mathcal{V}$ is uniformly bounded if $Diam(\mathcal{V})=\sup\{diam(V)\mid V\in\mathcal{V}\}$ is finite. Let $\mathcal{U}$ be a cover of $X$. The Lebesgue number of $\mathcal{U}$, denoted as $L(\mathcal{U})$, is the largest number $R$ such that for every subset $V$ of $X$ with $diam(V)\leq R$ there exists $U\in\mathcal{U}$ such that $V\subset U$. The multiplicity of $\mathcal{U}$, denoted as $m(\mathcal{U})$, is the largest number of elements of $\mathcal{U}$ with nonempty intersection. The $R$-multiplicity of $\mathcal{U}$, denoted as $m_{R}(\mathcal{U})$, is the smallest integer $n$ such that for any $x\in X$, the ball $B(x,R)$ intersects at most $n$ sets of $\mathcal{U}$.
Definition 2.1 ([Dra06]).
Let $(X,d)$ be a metric space. For any $R>0$, let $\textnormal{ad}_{X}(R)$ be the smallest number over $n\in\mathbb{N}$ for which there exists a uniformly bounded cover $\mathcal{U}=\{U_{i}\}_{i\in I}$ of $X$ with $L(\mathcal{U})>R$ and $m(\mathcal{U})\leq n+1$. We call $\textnormal{ad}_{X}:\mathbb{R}^{+}\rightarrow\mathbb{N}$ the asymptotic dimension function of $X$.
Remark 2.2.
It is straightforward to see that
  1. (1)
    The function $\textnormal{ad}_{X}$ is nondecreasing, and the asymptotic dimension of $X$ is defined as follows:
    $\textnormal{asdim}(X)=\lim\limits_{R\rightarrow\infty}\textnormal{ad}_{X}(R).$ (2.1)
  2. (2)
    $\textnormal{asdim}(X)\leq d$ if and only if $\textnormal{ad}_{X}$ is constant and $\textnormal{ad}_{X}\leq d$.
  3. (3)
    If $\textnormal{asdim}(X)=\infty$, then both sides of Equation 2.1 are infinite.
Analogous to the volume function, the growth type of the asymptotic dimension function is an essential concept. Recall from [Bel05] that for $f,g:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$,
  • we call $f$ is at most $g$, denoted as $f\preceq g$, if there exists $k\in\mathbb{N}$ such that $f(x)\leq kg(kx+k)+k$ for any $x>k$;
  • we call $f$ is equivalent to $g$, denoted as $f\approx g$, if both $f\preceq g$ and $g\preceq f$. Clearly, $\approx$ is an equivalence relation;
  • we define the growth type of $f$ to be the $\approx$-equivalence class of $f$;
  • we define the asymptotic dimension growth of $X$ to be the growth type of $\textnormal{ad}_{X}$.
Although the asymptotic dimension growth is not a coarse invariant like the finite asymptotic dimension, it is a quasi-isometric invariant by the following result. Hence, it is a group invariant as any two word metrics on a group are quasi-isometric.
Proposition 2.3 ([Bel05, Dra06]).
Let $X$ and $Y$ be two discrete metric spaces with bounded geometry. If $X$ and $Y$ are quasi-isometric, then $\textnormal{ad}_{X}\approx\textnormal{ad}_{Y}$. In particular, the asymptotic dimension growth is well-defined for finitely generated groups.
Below, we will introduce three equivalent definitions of asymptotic dimension growth. The first equivalent definition shown in [[ANWZ18], Lemma 2.6] links the previous and the following.
Lemma 2.4 ([ANWZ18]).
Let $(X,d)$ be a metric space. For any $R>0$, we define $\widetilde{\textnormal{ad}_{X}}(R)$ to be the smallest number over $n\in\mathbb{N}$ for which there exists a uniformly bounded cover $\mathcal{V}=\{V_{j}\}_{j\in J}$ with $m_{R}(\mathcal{V})\leq n+1$. Then $\textnormal{ad}_{X}\approx\widetilde{\textnormal{ad}_{X}}$.
Next, we give the second equivalent definition of the asymptotic dimension growth.
Proposition 2.5.
Let $(X,d)$ be a separable metric space. For any $R>0$, we define $\widehat{\textnormal{ad}_{X}}(R)$ to be the smallest number over $n\in\mathbb{N}$ for which there exist uniformly bounded, $R$-disjoint families $\mathcal{W}_{0},\cdots,\mathcal{W}_{n}$ of subsets of $X$ such that $\bigcup\limits_{l=0}^{n}\mathcal{W}_{l}$ is a cover of $X$. Then $\widetilde{\textnormal{ad}_{X}}\approx\widehat{\textnormal{ad}_{X}}$, hence $\textnormal{ad}_{X}\approx\widehat{\textnormal{ad}_{X}}$.
Proof.
Given $R>0$, $2R>R$. Suppose that there are uniformly bounded, $2R$-disjoint families $\mathcal{W}_{0},\cdots,\mathcal{W}_{\widehat{\textnormal{ad}_{X}}(2R)}$ of subsets of $X$ covering $X$. Put $\mathcal{W}=\bigcup\limits_{l}\mathcal{W}_{l}$, which is obviously a uniformly bounded cover. We claim that $m_{R}(\mathcal{W})\leq\widehat{\textnormal{ad}_{X}}(2R)$. By definition of $R$-multiplicity, for any $x\in X$, there are at most $\widehat{\textnormal{ad}_{X}}(2R)+1$ of sets of $\mathcal{W}$ having non-empty intersection with the closed ball $B(x,R)$. Otherwise, there must be two sets $W$ and $W^{\prime}$ from the same family $\mathcal{W}_{i}$ for some $i$ such that $W\cap B(x,R)\neq\varnothing$ and $W^{\prime}\cap B(x,R)\neq\varnothing$, since there are only $\widehat{\textnormal{ad}_{X}}(2R)+1$ different families. Then $d(W,W^{\prime})\leq 2R$, which contradicts that the family $\mathcal{W}_{i}$ is $2R$-disjoint. Hence, $m_{R}(\mathcal{W})\leq\widehat{\textnormal{ad}_{X}}(2R)$ as claimed, and thus yields $\widetilde{\textnormal{ad}_{X}}\preceq\widehat{\textnormal{ad}_{X}}$.
Conversely, given $R>0$, suppose $\mathcal{V}=\{V_{j}\}_{j\in J}$ is a unifromly bounded cover with $m_{R}(\mathcal{V})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$. Since $X$ is separable, $X$ is second countable, so a cover of $X$ must have a countable subcover. Thus we may assume that the cover $\mathcal{V}$ is countable. We first claim that there exist uniformly bounded, $R$-disjoint families $\mathcal{V}_{0},\cdots,\mathcal{V}_{\widetilde{\textnormal{ad}_{X}}(R)}$ of subsets of $X$ covering $X$. We prove the first claim in three steps.
Step 1. In the first step, we construct an irreducible cover of $X$. Let $\mathcal{U}=\{U_{i}\}_{i\in I}$ be a subcover of $\mathcal{V}$ such that
$V_{j}\bigg{\backslash}\left(\bigcup\limits_{V^{\prime}\in\mathcal{V}\backslash\{V_{j}\}}V^{\prime}\right)\neq\varnothing,\text{ for all }V_{j}\in\mathcal{V},$
which is an irreducible cover of $X$. As $m_{R}(\mathcal{U})\leq m_{R}(\mathcal{V})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$, $m_{R}(\mathcal{U})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$ ( here the equality can be obtained ).
Step 2. In this step, we build $\widetilde{\textnormal{ad}_{X}}(R)+1$ boxes. Since $m_{R}(\mathcal{U})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$ and the equality can be obtained, there exists $x\in X$ such that $B(x,R)$ intersects $\widetilde{\textnormal{ad}_{X}}(R)+1$ elements $U_{0},\cdots,U_{\widetilde{\textnormal{ad}_{X}}(R)}$ of $\mathcal{V}$. Thus $d(U_{i},U_{j})\leq 2R$, for all $i,j\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$ with $i\neq j$, and we put $U_{0},\cdots,U_{\widetilde{\textnormal{ad}_{X}}(R)}$ into $\widetilde{\textnormal{ad}_{X}}(R)+1$ different boxes.
Step 3. In the final step, we put the remaining sets of the cover $\mathcal{U}$ into these $\widetilde{\textnormal{ad}_{X}}(R)+1$ boxes. Set $\mathcal{U}\big{\backslash}\bigcup\limits_{i=0}^{\widetilde{\textnormal{ad}_{X}}(R)}U_{i}=\{U_{k}\}_{k\in I}$, and we denote by $\Box_{l}$ the $l$-th box. We next claim that for any $x\in U_{k}$, there exists $l\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$ such that
$d(x,B_{l})>R,\forall B_{l}\in\Box_{l}.$
Otherwise, there exists $y\in U_{k}$ such that for all $l\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$, there is a set $B_{l}\in\Box_{l}$ such that
$d(y,B_{l})\leq R,$
which implies that $B(y,R)$ intersects $\widetilde{\textnormal{ad}_{X}}(R)+2$ sets $U_{k},B_{0},\cdots,B_{\widetilde{\textnormal{ad}_{X}}(R)}$ of $\mathcal{U}$, which contradicts that $m_{R}(\mathcal{U})\leq\widetilde{\textnormal{ad}_{X}}(R)+1$. Hence, for any $x\in U_{k}$, $x$ must be in some box. Below, we divide $U_{k}$ into several pieces and put them into boxes one by one. Let
$U^{(0)}_{k}$ =$=\{x\in U_{k}\mid d(x,B_{0})>R,\text{ for all }B_{0}\in\Box_{0}\},$
$U^{(l)}_{k}$ =$=\{x\in U_{k}\big{\backslash}\bigcup\limits_{n=0}^{l-1}U^{(n)}_{k}\mid d(x,B_{l})>R,\text{ for all }B_{l}\in\Box_{l}\},l=1,\cdots,\widetilde{\textnormal{ad}_{X}}(R).$
Then we put $U^{(l)}_{k}$ into the box $\Box_{l}$ for $l=0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)$ in ascending order. If $U^{(l)}_{k}$ is the empty set, we consider the next box. Observe that $U^{(l)}_{k}$ is a subset of $U_{k}$ and $U_{k}=\bigsqcup\limits_{l=0}^{\widetilde{\textnormal{ad}_{X}}(R)}U^{(l)}_{k}$. Similarly, we can divide all elements of $\mathcal{U}\big{\backslash}\bigcup\limits_{j=0}^{\widetilde{\textnormal{ad}_{X}}(R)}U_{j}$ into $\widetilde{\textnormal{ad}_{X}}(R)+1$ boxes. Therefore, we obtain new boxes and for all $l\in\{0,\cdots,\widetilde{\textnormal{ad}_{X}}(R)\}$,
$d(B_{l},B^{\prime}_{l})>R,\text{ for every }B_{l}\neq B^{\prime}_{l}\text{ in }\Box_{l}.$
Let $\mathcal{W}_{l}=\{B_{l}\mid B_{l}\in\Box_{l}\}$, then each $\mathcal{W}_{l}$ is a uniformly bounded, $R$-disjoint family and $X=\bigcup\limits_{l=0}^{\widetilde{\textnormal{ad}_{X}}(R)}\mathcal{W}_{l}$. By definition, it is obvious that $\widehat{\textnormal{ad}_{X}}(R)\leq\widetilde{\textnormal{ad}_{X}}(R)$, which yields $\widehat{\textnormal{ad}_{X}}\preceq\widetilde{\textnormal{ad}_{X}}$.
By the transitivity of the equivalence relation, we obtain that $\textnormal{ad}_{X}\approx\widehat{\textnormal{ad}_{X}}$ by Lemma 2.4. ∎
Next, we provide a couple of typical examples of groups and spaces with asymptotic dimension growth. We mainly consider four classes of growth types: exponential growth, subexponential growth, polynomial growth, and slow growth. Say that a monotone increasing function $f:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ has subexponential growth if for any $b>1$,
$\lim\limits_{t\rightarrow\infty}\frac{f(t)}{b^{t}}=0.$
Observe that if $f$ grows polynomially, it obviously grows subexponentially, since polynomial functions grow slower than exponential functions. Say that $X$ has polynomial (resp. subexponential, exponential) asymptotic dimension growth, if there exists a polynomial (resp. subexponential, exponential) function $f$ such that $\textnormal{ad}_{X}\preceq f$. Recall from [Dra00] that $X$ has slow asymptotic dimension growth, if there exists a linear function $f$ such that $\textnormal{ad}_{X}\preceq f$.
Like the volume growth, a finitely generated group has at most exponential asymptotic dimension growth.
Example 2.6 ([Dra06]).
For a finitely generated group $\Gamma$ with the word metric, there exists $\alpha>0$ such that $ad_{\Gamma}(R)\preceq e^{\alpha R}$.
In [Oza12], Ozawa showed that a metric space with subexponential asymptotic dimension growth has property A. Therefore, the following two examples have property A.
Example 2.7 ([ANWZ18]).
Let $X$ be a geodesic coarse median space with finite rank and at most exponential volume growth, then $X$ has subexponential asymptotic dimension growth.
Example 2.8 ([Dra06]).
Let $N$ be a finitely generated nilpotent group and $G$ be a finitely generated group with finite asymptotic dimension. Then the wreath product $N\wr G$ has polynomial asymptotic dimension growth. In particular, $\mathbb{Z}\wr\mathbb{Z}$ has polynomial asymptotic dimension growth.
The following example tells us that given a strictly increasing function $f$, we can construct a topological space of asymptotic dimension function $f$. This provides an approach to build spaces with slow asymptotic dimension growth when $f$ grows slowly.
Example 2.9 ([CFY09]).
Let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a strictly increasing function. Let $G$ be the direct product given by $G=\bigoplus\limits_{n\in\mathbb{N}}f(n)\mathbb{Z}$, and endow $G$ with the direct sum metric. Then $G$ is a discrete metric space of bounded geometry with asymptotic dimension function $f$.
2.2. The case of coarse spaces
In this subsection, we specialize in the case of coarse spaces. In the case of a coarse space for which the coarse structure is induced from the metric, we observe that $R$-disjointness and uniformly boundedness can be converted into $E$-separatedness and $E$-boundedness in coarse geometry. Further, we show that statements about finite asymptotic dimensions on metric spaces and coarse spaces are equivalent. Moreover, we will define the asymptotic dimension function for such coarse spaces, and prove that the function is the same as the function defined in Proposition 2.5. In the rest of this paper, we will study separable coarse spaces.
Definition 2.10 ([Roe03]).
Let $X$ be a set, and let $\mathcal{E}$ be a collection of subsets of $X\times X$. We say that $\mathcal{E}$ is a coarse structure on $X$ if it satisfies the following axioms:
  • The diagonal =$\bigtriangleup=\{(x,x)\mid x\in X\}$ is a member of $\mathcal{E}$;
  • $E\in\mathcal{E}$ and $F\subset E$ implies $F\in\mathcal{E}$;
  • $E,F\in\mathcal{E}$ implies $E\cup F\in\mathcal{E}$;
  • $E\in\mathcal{E}$ implies $E^{-1}:=\{(y,x)\mid(x,y)\in E\}\in\mathcal{E}$;
  • $E,F\in\mathcal{E}$ implies $E\circ F:=\{(x,y)\mid\exists z,s.t.(x,z)\in E\text{ and }(z,y)\in F\}\in\mathcal{E}$;
The pair $(X,\mathcal{E})$ is called a coarse space, and the elements of $\mathcal{E}$ are called controlled sets or entourages.
Remark 2.11.
Below, we consider the special case of the coarse structure coming from a metric. Let $(X,d)$ be a discrete metric space. For $R>0$, we define a controlled tube
$E_{R}:=\{(x,y)\in X\times X\mid d(x,y)\leq R\}.$
Then the controlled sets are defined as subsets $E$ of $X\times X$ contained in some controlled tube $E_{R}$ for some $R$, which form a coarse structure on $X$.
From the perspective of coarse geometry, we can correspond $R$-disjointness and uniformly boundedness to $E$-separatedness and $E$-boundedness in the following definition.
Definition 2.12.
[Roe03] Let $X$ be a coarse space, and let $E$ be a controlled set for $X$. A family $\mathcal{U}=\{U_{i}\}_{i\in I}$ subsets of $X$ is:
  1. (1)
    $E$-separated if $U_{i}\times U_{j}\cap E=\varnothing$ when $i\neq j$;
  2. (2)
    $E$-bounded if $U_{i}\times U_{i}\subset E$ for each $i\in I$.
For the finite asymptotic dimension, the statements on the coarse space and the metric space are equivalent in the above special case.
Lemma 2.13.
Let $X$ be a coarse space. Then the following two conditions are equivalent:
  1. (1)
    for any controlled set $E$, there exist a controlled set $F$ and a cover $\mathcal{U}=\{U_{i}\}_{i\in I}$ of $X$ such that $\mathcal{U}$ is $F$-bounded and such that $\mathcal{U}$ admits a decomposition
    $\mathcal{U}=\mathcal{U}_{i}\sqcup\cdots\sqcup\mathcal{U}_{n}$
    such that each $\mathcal{U}_{i}$ is $E$-separated.
  2. (2)
    for any $R>0$, there exist uniformly bounded, $R$-disjoint families $\mathcal{U}_{0},\cdots,\mathcal{U}_{n}$ of subsets of $X$ such that $\bigcup\limits_{i=0}^{n}\mathcal{U}_{i}$ covers $X$.
Remark 2.14.
We can complete the proof by the method analogous to the next lemma.
In the case of coarse spaces, we give an alternative definition of the asymptotic dimension function equal to the function defined in Proposition 2.5. This definition plays an important role in the proof of Theorem 4.12 in Section 4.
Lemma 2.15.
Let $X$ be a separable metric space. For any $R>0$, we define $\widecheck{\textnormal{ad}_{X}}(R)$ to be the smallest number over all $m\in\mathbb{N}$ with the following property: for the controlled set $E_{R}$, there exist a controlled set $F$ and a cover $\mathcal{U}=\{U_{i}\}_{i\in I}$ of $X$ such that $\mathcal{U}$ is $F$-bounded and such that $\mathcal{U}$ admits a decomposition
$\mathcal{U}=\mathcal{U}_{0}\sqcup\cdots\sqcup\mathcal{U}_{m}$
such that each $\mathcal{U}_{i}$ is $E_{R}$-separated. Then $\widecheck{\textnormal{ad}_{X}}=\widehat{\textnormal{ad}_{X}}$, hence ${\textnormal{ad}_{X}}\approx\widecheck{\textnormal{ad}_{X}}$.
Proof.
Given $R>0$. Suppose that $\mathcal{U}_{0},\cdots,\mathcal{U}_{\widehat{\textnormal{ad}_{X}}(R)}$ are uniformly bounded, $R$-disjoint families of subsets of $X$ covering $X$. Since each family $\mathcal{U}_{i}$ is uniformly bounded, the cover $\mathcal{U}=\bigsqcup\limits_{i=0}^{\widehat{\textnormal{ad}_{X}}(R)}\mathcal{U}_{i}$ is uniformly bounded. It follows that there exists a constant $C>0$ such that for each $U\in\mathcal{U}$, $\mathrm{diam}(U)\leq C$, which implies that each set $U\times U$ is contained in a controlled set $E_{C}$, thus the cover $\mathcal{U}$ is $E_{C}$-bounded. As each family $\mathcal{U}_{i}$ is $R$-disjoint, then for every $U\neq U^{\prime}$ in $\mathcal{U}_{i}$, $d(U,U^{\prime})>R$, which implies that $(U\times U^{\prime})\cap E_{R}=\varnothing$, so each $\mathcal{U}_{i}$ is $E_{R}$-separated. Hence, $\widecheck{\textnormal{ad}_{X}}(R)\leq\widehat{\textnormal{ad}_{X}}(R)$, which yields $\widecheck{\textnormal{ad}_{X}}\leq\widehat{\textnormal{ad}_{X}}$.
Conversely, given $R>0$, for the controlled set $E_{R}$, there exist a controlled set $F$ and a cover $\mathcal{U}=\{U_{i}\}_{i\in I}$ of $X$ such that $\mathcal{U}$ is $F$-bounded and such that $\mathcal{U}$ admits a decomposition
$\mathcal{U}=\mathcal{U}_{0}\sqcup\cdots\sqcup\mathcal{U}_{\widecheck{\textnormal{ad}_{X}}(R)}$
such that each $\mathcal{U}_{i}$ is $E_{R}$-separated. Since $F$ is a controlled set, $F\subseteq E_{C}$ for some $C>0$. As $\mathcal{U}$ is $F$-bounded, each set $U_{i}\times U_{i}$ is contained in $E_{C}$, which implies that $diam(U_{i})\leq C$, so $\mathcal{U}$ is a uniformly bounded cover. Since each $\mathcal{U}_{i}$ is $E_{R}$-separated, for every $U\neq U^{\prime}$ in $\mathcal{U}_{i}$, $(U\times U^{\prime})\cap E_{R}=\varnothing$, thus $d(U,U^{\prime})>R$, which implies that each $\mathcal{U}_{i}$ is $R$-disjoint. Therefore, $\widehat{\textnormal{ad}_{X}}(R)\leq\widecheck{\textnormal{ad}_{X}}(R)$, which yields $\widehat{\textnormal{ad}_{X}}\leq\widecheck{\textnormal{ad}_{X}}$. Hence, $\widehat{\textnormal{ad}_{X}}=\widecheck{\textnormal{ad}_{X}}$, which implies that $\textnormal{ad}_{X}\approx\widecheck{\textnormal{ad}_{X}}$ by Proposition 2.5. ∎
3. Dynamic asymptotic dimension growth for group actions
In this section, we will introduce the dynamic asymptotic dimension function and the dynamic asymptotic dimension growth for group actions on compact spaces. For a nondecreasing function $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$, we define the $h$-BLR action, which is inspired by the work of Bartels, Reich, and Lück [BLR08, Bar16]. Furthermore, we will show that an $h$-BLR action has dynamic asymptotic dimension function at most $h$. Also, when the function $h$ grows slower than $\sqrt{x}$, we will obtain an equivariant condition of the BLR condition, which is similar to the definition of asymptotic dimension growth for groups. Moreover, it bridges the dynamic asymptotic dimension growth for group actions with the asymptotic dimension growth for groups in the sense of coarse geometry in Section 4, which allows us to construct several exciting examples of dynamic asymptotic dimension growth for group actions.
3.1. The case of group actions on compact spaces
In this subsection, we will introduce the dynamic asymptotic dimension growth for group actions on compact spaces, which is the dynamical version of asymptotic dimension growth. Moreover, we will give some interesting examples.
To introduce our primary definition, we first recall the notion of length functions on discrete groups.
Definition 3.1.
Let $\Gamma$ be a discrete group. A proper length function on $\Gamma$ is a function $\ell:\Gamma\rightarrow[0,\infty)$ satisfying:
  1. (1)
    $\ell(\gamma)=0$ if and only if $\gamma=e$;
  2. (2)
    $\ell(\gamma_{1}+\gamma_{2})\leq\ell(\gamma_{1})+\ell(\gamma_{2})$, for all $\gamma_{1},\gamma_{2}\in\Gamma$;
  3. (3)
    $\ell(\gamma^{-1})=\ell(\gamma)$ for all $\gamma\in\Gamma$;
  4. (4)
    $\{\gamma\in\Gamma\mid\ell(\gamma)\leq R\}$ is finite for all $R\geq 0$.
Using the concept of proper length functions on discrete groups, we can define the dynamic asymptotic dimension growth for group actions, which is a generalization of the finite dynamic asymptotic dimension for group actions.
Definition 3.2.
Let $\Gamma$ be a discrete group, and let $X$ be a compact Hausdorff space. Let $\Gamma\curvearrowright X$ be an action by homeomorphisms, and let $\ell_{\Gamma}:\Gamma\rightarrow[0,\infty)$ be a proper length function. For any $R>0$, let $\textnormal{dad}_{\Gamma\curvearrowright X}(R)$ be the smallest number over $m\in\mathbb{N}$ with the following property: for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$, there is an open cover $\{U_{0},\cdots,U_{m}\}$ of $X$ such that for each $i\in\{0,\cdots,m\}$, the set
[…]
is finite. We call $\textnormal{dad}_{\Gamma\curvearrowright X}:\mathbb{R}^{+}\rightarrow\mathbb{N}$ the dynamic asymptotic dimension function of $\Gamma\curvearrowright X$, and call the growth type of the function $\textnormal{dad}_{\Gamma\curvearrowright X}$ the dynamic asymptotic dimension growth of $\Gamma\curvearrowright X$.
Remark 3.3.
  1. (1)
    The function $\textnormal{dad}_{\Gamma\curvearrowright X}$ is non-decreasing, and the dynamic asymptotic dimension of the action $\Gamma\curvearrowright X$ is defined as follows:
    $\textnormal{dasdim}(\Gamma\curvearrowright X)=\lim\limits_{R\rightarrow\infty}\textnormal{dad}_{\Gamma\curvearrowright X}(R).$ (3.1)
  2. (2)
    $\textnormal{dasdim}(\Gamma\curvearrowright X)\leq d$ for some finite $d$ if and only if $\textnormal{dad}_{\Gamma\curvearrowright X}$ is a constant function and $\textnormal{dad}_{\Gamma\curvearrowright X}\leq d$.
  3. (3)
    If $\textnormal{dasdim}(\Gamma\curvearrowright X)=\infty$, then both sides of Equation 3.1 are infinite.
  4. (4)
    $E=E^{-1}$ and $e\in E$ since $\ell$ is a proper length function.
Combining theorems in Section 4 and examples of groups with subexponential asymptotic dimension growth, we can provide several examples of group actions with subexponential dynamic asymptotic dimension growth.
Example 3.4.
Let $N$ be a finitely generated nilpotent group, and $\Gamma$ be a finitely generated group with finite asymptotic dimension. Then the wreath product $N\wr\Gamma$ admits a free, minimal action on the Cantor set which has polynomial dynamic asymptotic dimension growth. In particular, $\mathbb{Z}\wr\mathbb{Z}$ admits a free, minimal action on the Cantor set which has polynomial dynamic asymptotic dimension growth.
Example 3.5.
Let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a strictly increasing function with $f\preceq x$. Let $G$ be the direct product given by $G=\bigoplus\limits_{n\in\mathbb{N}}f(n)\mathbb{Z}$, and endow $G$ with the direct sum metric. Then $G$ admits a free, minimal action on the Cantor set which has slow dynamic asymptotic dimension growth.
3.2. The BLR condtions
In this subsection, we will introduce the BLR condition for a group action derived from the work of Bartels, Reich, and Lück on the Farrell-Jones conjecture [BLR08, Bar16]. There are several benefits to considering this notion. Firstly, it is similar to the definition of the amenability of a group action. Secondly, it is equivalent to the asymptotic dimension growth for a group under certain conditions. Finally, it is closely related to the dynamic asymptotic dimension growth for group actions.
To define the $h$-BLR action, we need a couple of definitions.
Definition 3.6.
Let $V$ be a countable set. The space of probability measures on $V$ is
$\textnormal{Prob}(V):=\{\mu\in\ell^{1}(V)\mid\mu(v)\geq 0,\forall v\in V\text{and }\|\mu\|_{1}=1\},$
equipped with the metric
$d\left(\sum\limits_{v\in V}b_{v}v,\sum\limits_{v\in V}c_{v}v\right):=\sum\limits_{v\in V}|b_{v}-c_{v}|.$
coming from $\ell^{1}$-norm.
Remark 3.7.
Elements of $\textnormal{Prob}(V)$ can be written as $\mu=\sum\limits_{v\in V}t_{v}v$, where $\mu(v)=t_{v}$ in $[0,1]$ and $\sum\limits_{v\in V}t_{v}=1$. Write $\textnormal{Prob}_{n}(V)$ for the subset of $\textnormal{Prob}(V)$ consisting of measures supported on at most $n+1$ points. A simplicial complex is a subspace $C$ of some $\bigcup\limits_{n=0}^{\infty}\textnormal{Prob}_{n}(V)$ such that $V\subset C$ and
$\left\{\sum\limits_{v\in V}t_{v}v\in\bigcup\limits_{n=0}^{\infty}\textnormal{Prob}_{n}(V)\mid t_{v}\neq 0\text{ for }v\in\textnormal{supp }(\mu)\right\}\in C.$
Clearly, $V$ is the vertex set of $C$.
Definition 3.8 ([GWY17]).
Let $X$ be a topological space, and let $Y$ be a metric space. Assume that $\Gamma$ is a discrete group acting on $X$ by homeomorphisms, and on $Y$ by isometries. Let $E$ be a subset of $\Gamma$ and $\varepsilon>0$. Say that a map $f:X\rightarrow Y$ is $(E,\varepsilon)$-equivariant if
$\sup\limits_{x\in X}d_{Y}(f(gx),gf(x))<\varepsilon,\forall g\in E$
To describe the following lemma, we review the notion of a simplicial action. Let $\Gamma$ be a discrete group, and let $C$ be a simplicial complex with the vertex set $V$. A simplicial action $g$ of $\Gamma$ on $C$ is an action that is induced from some action of $\Gamma$ on $V$ via the formula
$g\left(\sum\limits_{v\in V}t_{v}v\right)=\sum\limits_{v\in V}t_{v}(gv).$
Clearly, a simplicial action is isometric. The next technical lemma will be used in the proof of Proposition 3.12 and Theorems 3.14 and 4.17.
Lemma 3.9 ([GWY17]).
Let $f:X\rightarrow C$ be a continuous $(E,\varepsilon)$-equivariant map as in the above definition, where $X$ is compact and $E$ is finite. Then there exists a finite subset $S$ of the vertex set of $C$ and an $(E,\varepsilon)$-equivariant map $f^{\prime}:X\rightarrow C$ such that $f^{\prime}(X)\subseteq\textnormal{Prob}(S)\cap C$.
Definition 3.10 ([GWY17]).
Let $C$ be a simplicial complex equipped with a simplicial action of a discrete group $\Gamma$. Let $\mathcal{F}$ be a family of subgroups of $\Gamma$ satisfying the following conditions:
  1. (1)
    $\mathcal{F}$ is closed under conjugation;
  2. (2)
    $\mathcal{F}$ is closed under taking subgroups;
  3. (3)
    $\mathcal{F}$ is closed under taking finite index supergroups.
Then $C$ is called a $(\Gamma,\mathcal{F})$-complex if the stabilizer of every vertex in $C$ is an element of $\mathcal{F}$.
The $d$-BLR condition was first defined in Bartels’ paper [Bar16], then discussed in the paper [GWY17] to connect with the finite asymptotic dimension. Below, we will generalize it to the growth case.
Definition 3.11.
Let $\Gamma\curvearrowright X$ be an action with $X$ compact, and let $\mathcal{F}$ be a family of subgroups of $\Gamma$. Let $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function. We say that the action is $h$-BLR for $\mathcal{F}$ if for any $R>0$, for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$ and any $\varepsilon>0$, there exist a $(\Gamma,\mathcal{F})$-complex $C$ of dimension at most $h(R)$ and a continuous $(E,\varepsilon)$-equivariant map $f:X\rightarrow C$.
If $h\preceq x^{\alpha}(0<\alpha<\frac{1}{2})$, then we get an equivariant version of the $h$-BLR condition, which is associated with the asymptotic dimension growth of groups.
Proposition 3.12.
Let $\Gamma\curvearrowright X$ be an action with $X$ compact, and let $\mathcal{F}$ be a family of subgroups of $\Gamma$. Let $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function with $h\preceq x^{\alpha}(0<\alpha<\frac{1}{2})$. Then the following are equivalent.
  1. (1)
    The action is $h$-$BLR$ for $\mathcal{F}$.
  2. (2)
    For any $R>0$, there exists a nondecreasing function $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ such that for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$, there exists a $\Gamma$ equivaraint open cover $\mathcal{U}$ of $X\times\Gamma$ such that the following hold.
    1. (a)
      For each $U\in\mathcal{U}$ and $\gamma\in\Gamma$, $\gamma U=U$ or $\gamma U\cap U=\varnothing$.
    2. (b)
      For each $U\in\mathcal{U}$, $\{\gamma\in\Gamma\mid\gamma U=U\}\in\mathcal{F}$.
    3. (c)
      $m(\mathcal{U})\leq h(R)+1$.
    4. (d)
      $\mathcal{U}/\Gamma$ is finite.
    5. (e)
      For each $\gamma\in\Gamma$ and $x\in X$, there exists $U\in\mathcal{U}$ such that $\{x\}\times\gamma E\subset U$.
Remark 3.13.
This proposition can be proved in a similar way as shown in [Proposition 4.5, [GWY17]]. In the proof of $(1)$ implying $(2)$, there is no extra restrictions on the growth rate of $h$, while in the case of $(2)$ implying $(1)$, the function $h$ must grow slower than $\sqrt{x}$.
The following theorem shows that an $h$-BLR action has dynamic asymptotic dimension function at most $h$. Under the assumption of subexponential growth of $h$, Theorem 4.17 will give a converse in coarse geometry in the next section.
Theorem 3.14.
Let $\Gamma\curvearrowright X$ be an action with $X$ compact, and let $\mathcal{F}$ be the family of finite subgroups of $\Gamma$. Let $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function. If $\Gamma\curvearrowright X$ is an $h$-BLR action for $\mathcal{F}$, then it has dynamic asymptotic dimension function at most $h$.
Proof.
First, we construct a desired open cover of $X$. Assume that $\Gamma\curvearrowright X$ is an $h$-$BLR$ action for $\mathcal{F}$. Given $R>0$, there exists a nondecreasing function $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ such that for the finite subset $E=\{\gamma\in\Gamma\mid\ell(\gamma)<R\}$ of $\Gamma$ and $\varepsilon=\frac{1}{3}10^{-h(R)}$, there exist a $(\Gamma,\mathcal{F})$-complex $C$ of dimension at most $h(R)$ and a continuous $(E,\varepsilon)$-equivariant map $f:X\rightarrow C$. By Lemma 3.9, we may assume that there exists a finite subset $S$ of the vertex set of $C$ such that $f(X)\subseteq C\cap\textnormal{Prob}(S)$. Define
$F:=\{g\in\Gamma\mid gS\cap S\neq\varnothing\}.$
Since $S$ is finite and the stabilizer of each vertex in $S$ is finite, $F$ is a finite subset of $\Gamma$. Observe that $C$ is a $(\Gamma,\mathcal{F})$-complex $C$ of dimension at most $h(R)$, there exists an open cover $\{V_{0},\cdots,V_{h(R)}\}$ of $C$ as in [[GWY17], Lemma 4.1]. For each $i\in\{0,\cdots,h(R)\}$, and each $i$-simplex $\Delta$, we define $U_{i}:=f^{-1}(V_{i})$, $U_{i\Delta}:=f^{-1}(V_{i\Delta})$. Therefore, $\{U_{0},\cdots,U_{h(R)}\}$ is an open cover of $X$, and $U_{i}=\bigsqcup\limits_{\Delta\text{ an i-simplex }}U_{i\Delta}$. Indeed, $f(X)\subseteq C\cap\textnormal{Prob}(S)\subseteq\bigcup\limits_{i=0}^{h(R)}V_{i}$ implies that $X\subseteq f^{-1}(\bigcup\limits_{i=0}^{h(R)}V_{i})=\bigcup\limits_{i=0}^{h(R)}f^{-1}(V_{i})=\bigcup\limits_{i=0}^{h(R)}U_{i}$. By [[GWY17], Lemma 4.1], we know that $V_{i\Delta}$ and $V_{i\Delta^{\prime}}$ are distinct for different $i$-simplices, which implies that $f^{-1}(V_{i\Delta})$ and $f^{-1}(V_{i\Delta^{\prime}})$ are distinct, hence $U_{i\Delta}$ and $U_{i\Delta^{\prime}}$ are distinct.
Next, we claim that the following statement holds for any $i$ and $i$-simplex $\Delta$:
$x\in U_{i\Delta},gx\in U_{i},g\in E\Rightarrow gx\in U_{ig\Delta}.$
Indeed, we see that $gx\in U_{ig\Delta^{\prime}}$ for some $i$-simplex $\Delta^{\prime}$. Since $f$ is $(E,\varepsilon)$-equivariant, we have
$d(f(gx),gf(x))<\varepsilon.$
As $x\in U_{i\Delta}$, $f(x)\in V_{i\Delta}$, which implies that $gf(x)\in V_{ig\Delta}$. Clearly, $gx\in f^{-1}(V_{i\Delta^{\prime}})$ implies $f(gx)\in V_{i\Delta^{\prime}}$. Observe that $d(V_{i\Delta^{\prime}},V_{ig\Delta})\geq\frac{1}{3}10^{-h(R)}$ when $\Delta^{\prime}\neq g\Delta$, but $d(f(gx),gf(x))<\frac{1}{3}10^{-h(R)}$, which forces $\Delta^{\prime}=g\Delta$. Hence, $gx\in U_{ig\Delta}$ as claimed.
Finally, we show that the action $\Gamma\curvearrowright X$ has dynamic asymptotic dimension function at most $h$. Suppose that $x\in U_{i}$ for some $i$ and $g=g_{n}\cdots g_{1}$ with $g_{k}\in E$ and $g_{k}\cdots g_{1}x\in U_{i}$, $\forall k\in\{1,\cdots,n\}$. If $x\in U_{i\Delta}$ for some $i$-simplex $\Delta$, then $gx\in U_{ig\Delta}$ by using the claim repeatedly, thus $f(gx)\in V_{ig\Delta}\cap f(X)\subseteq C\cap\textnormal{Prob}(S)$. Observe that $f(gx)$ and $f(x)$ are supported on vertices on $S$, which forces $gS\cap S\neq\varnothing$; thus $g$ belongs to the finite set $F$ By Definition 3.2, $\textnormal{dad}_{\Gamma\curvearrowright X}(R)\leq h(R)$, thus $\textnormal{dad}_{\Gamma\curvearrowright X}\leq h$. ∎
4. Dynamic asymptotic dimension growth for groupoids
In this section, we formulate the definition of dynamic asymptotic dimension growth for groupoids, which generalizes our notion for the case of group actions. Moreover, we will show that our notion is closely related to coarse geometry, the asymptotic dimension growth for countable discrete groups, and the amenability of groupoids.
Recall from [Ren80] that a groupoid consists of a set $\mathcal{G}$ of morphisms, a subset $\mathcal{G}^{(0)}\subseteq\mathcal{G}$ of obejects (called the unit space), together with source and range maps $s,r:\mathcal{G}\rightarrow\mathcal{G}^{(0)}$, a composition map
$\mathcal{G}^{(2)}:=\{(g_{1},g_{2})\in\mathcal{G}\times\mathcal{G}:s(g_{1})=r(g_{2})\}\ni(g_{1},g_{2})\mapsto g_{1}g_{2}\in\mathcal{G},$
and an inverse map $g\mapsto g^{-1}$ satisfying a couple of axioms. A locally compact Hausdorrf groupoid is equipped with a locally compact and Hausdorff topology such that structure maps are continuous. It is called étale if the range map $r$ (hence the source map $s$) is a local homeomorphism, i.e., for any $g\in\mathcal{G}$, there exists a neighborhood $U$ of $g$ such that $r(U)$ is open and $r|_{U}$ is a homeomorphism. From now on, we assume that groupoids are always locally compact and Hausdorff.
4.1. The case of étale groupoids
We now focus on a class of étale groupoids, where it is easier to find conditions to derive amenability. Using the notion of length functions on étale groupoids, we may define the dynamic asymptotic dimension growth for étale groupoids. Furthermore, we will show that this notion is a generalization of the dynamic asymptotic dimension growth for group actions in this subsection.
The notion of length functions on étale groupoids was first appeared in [OY19], and later generalized by Ma and Wu in their paper [MW20].
Definition 4.1 ([OY19, MW20]).
Let $\mathcal{G}$ be an étale groupoid. A length function on $\mathcal{G}$ is a function $\ell:\mathcal{G}\rightarrow[0,\infty)$ satisfying
  1. (1)
    $\ell(x)=0$ if and only if $x\in\mathcal{G}^{(0)}$;
  2. (2)
    $\ell(x)=\ell(x^{-1})$ for all $x\in\mathcal{G}$;
  3. (3)
    $\ell(xy)\leq\ell(x)+\ell(y)$, if $x$ and $y$ are composable in $\mathcal{G}$.
Furthermore, an étale groupoid enjoys the following properties.
Definition 4.2 ([MW20]).
Let $\mathcal{G}$ be an étale groupoid. Let $\ell:\mathcal{G}\rightarrow[0,\infty)$ be a length function on $\mathcal{G}$. For any subset $K\subseteq\mathcal{G}$, we write
¯$\bar{\ell}(K)=\sup\limits_{x\in K}\ell(x).$
We say $\ell$ is
  1. (1)
    proper if for any $K\subseteq\mathcal{G}\backslash\mathcal{G}^{(0)}$, ¯$\bar{\ell}(K)<\infty$ implies $K$ is precompact;
  2. (2)
    controlled if for any $K\subseteq\mathcal{G}$, $K$ is precompact implies ¯$\bar{\ell}(K)<\infty$;
  3. (3)
    coarse if it is both proper and controlled;
  4. (4)
    continuous if it is continuous function with regard to the topology of $\mathcal{G}$.
Remark 4.3.
When the unit space $\mathcal{G}^{(0)}$ is compact, the item (1) can be rewritten as: $\ell$ is proper if for any $K\subseteq\mathcal{G}$, ¯$\bar{\ell}(K)<\infty$ implies $K$ is precompact. It is straightforward to see that a continuous length function is controlled, hence the item (4) implies the item (2). Recall from [MW20] that two length functions $\ell_{1},\ell_{2}$ are said to be coarsely equivalent if for any $r>0$, we have
$\sup\{\ell_{1}(x):\ell_{2}(x)\leq r\}<\infty\text{ and }\sup\{\ell_{2}(x):\ell_{1}(x)\leq r\}<\infty.$
To obtain the amenability of a groupoid, we are mainly concerned about the coarse continuous length function. In the work of [MW20], Ma and Wu demonstrated the existence of coarse continuous length functions in the following case.
Theorem 4.4 ([MW20]).
Up to coarse equivalence, any $\sigma$-compact locally compact Hausdorff étale groupoid has a unique coarse continuous length function.
Here is a typical example of the above theorem.
Example 4.5 ([MW20]).
Let $\Gamma$ be a countable discrete group, and let $X$ be a $\sigma$-compact, locally compact Hausdorff space. The action $\Gamma\curvearrowright X$ by homeomorphisms induces an associated transformation groupoid $X\rtimes\Gamma$ as follows:
  • $X\rtimes\Gamma:=\{(\gamma x,\gamma,x)\mid x\in X,\gamma\in\Gamma\}$;
  • the unit space ${(X\rtimes\Gamma)}^{(0)}$ is $\{(x,e,x)\mid x\in X\}$, which is identified with $X$;
  • source and range maps: $s(\gamma x,\gamma,x)=x$, $r(\gamma x,\gamma,x)=\gamma x$;
  • composition: $(\gamma\gamma^{\prime}x,\gamma,\gamma^{\prime}x)(\gamma^{\prime}x,\gamma^{\prime},x)=(\gamma\gamma^{\prime}x,\gamma\gamma^{\prime},x)$;
  • inverse: $(\gamma x,\gamma,x)^{-1}=(x,\gamma^{-1},\gamma x)$.
The topology coming from $X\times\Gamma$, then $X\rtimes\Gamma$ is a $\sigma$-compact, locally compact étale groupoid. Let $\ell_{\Gamma}:\Gamma\rightarrow[0,\infty)$ be a proper length function on $\Gamma$, and let $g:X\rightarrow[0,\infty)$ be a continuous proper function on $X$. We define
$\ell_{X\rtimes\Gamma}:X\rtimes\Gamma\rightarrow[0,\infty),$
$\qquad\qquad\qquad\qquad\qquad\qquad(\gamma x,\gamma,x)\mapsto\ell_{\Gamma}(\gamma)(1+\max\{g(\gamma x),g(x)\}).$
Clearly, it is a coarse continuous length function on $X\rtimes\Gamma$. If $X$ is compact, we may simply choose $g=0$, and thus
$\ell_{X\rtimes\Gamma}(\gamma x,\gamma,x)=\ell_{\Gamma}(\gamma),\text{ for any }(\gamma x,\gamma,x)\in X\rtimes\Gamma.$
We are mainly interested in étale groupoids with compact unit spaces, and we now define the dynamic asymptotic dimension growth for such groupoids.
Definition 4.6.
Let $\mathcal{G}$ be an étale groupoid with compact unit space, and let $\ell:\mathcal{G}\rightarrow[0,\infty)$ be a coarse continuous length function. For any $R>0$, let $\textnormal{dad}_{\mathcal{G}}(R)$ be the smallest number over all $m\in\mathbb{N}$ with the following property: for the open precompact subset $K={\{z\in\mathcal{G}\mid\ell(z)<R\}}$ of $\mathcal{G}$, there exists an open cover $\{U_{0},\cdots,U_{m}\}$ of $\mathcal{G}^{(0)}$ such that for each $i$, the set $\{g\in K\mid s(g),r(g)\in U_{i}\}$ is contained in a precompact subgroupoid of $\mathcal{G}$. We call $\textnormal{dad}_{\mathcal{G}}:\mathbb{R}^{+}\rightarrow\mathbb{N}$ the dynamic asymptotic dimension function of $\mathcal{G}$, and call the growth type of the function $\textnormal{dad}_{\mathcal{G}}$ the dynamic asymptotic dimension growth of $\mathcal{G}$.
Remark 4.7.
Clearly, the function $\textnormal{dad}_{\mathcal{G}}$ is non-decreasing, and the dynamic asymptotic dimension of $\mathcal{G}$ is defined as follows:
$\textnormal{dasdim}(\mathcal{G})=\lim\limits_{R\rightarrow\infty}\textnormal{dad}_{\mathcal{G}}(R).$ (4.1)
  1. (1)
    $\textnormal{dasdim}(\mathcal{G})\leq d$ for some finite $d$ if and only if $\textnormal{dad}_{\mathcal{G}}$ is a constant function and $\textnormal{dad}_{\mathcal{G}}\leq d$.
  2. (2)
    If $\textnormal{dasdim}(\mathcal{G})=\infty$, then both sides of Equation 4.1 are infinite.
  3. (3)
    For each $i$, the subgroupoid $\mathcal{G}_{i}$ generated by the set
    $\{g\in K\mid s(g),r(g)\in U_{i}\}$
    is open precompact by the next lemma.
The following lemma will be used in the proof of a key lemma in Section 5.
Lemma 4.8 ([GWY17]).
Let $\mathcal{G}$ be an étale groupoid.
  1. (1)
    If $K,H$ are open (resp. compact, precompact) subsets of $\mathcal{G}$, then
    $K\cdot H:=\{kh\in\mathcal{G}\mid k\in K,h\in H,s(k)=r(h)\}$
    is open (resp. compact, precompact).
  2. (2)
    If $K\subseteq\mathcal{G}$ is open, then the subgroupoid of $\mathcal{G}$ generated by $K$ is also open, and is itself an étale groupoid.
To show that Definition 4.6 generalizes the dynamic asymptotic dimension growth for group actions in the sense of Definition 3.2, we consider the transformation groupoids induced by group actions on compact space $X$ in Example 4.5.
Lemma 4.9.
Let $X$ be a compact Hausdorff space, and let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function. An action $\Gamma\curvearrowright X$ has dynamic asymptotic dimension function $f$ if and only if the corresponding transformation groupoid $\mathcal{G}=X\rtimes\Gamma$ has dynamic asymptotic dimesnion function $f$.
Proof.
We may assume that an action $\Gamma\curvearrowright X$ has dynamic asymptotic dimension function $f$, and its associated transformation groupoid has dynamic asymptotic dimension funnction $\textnormal{dad}_{\mathcal{G}}$. It suffices to show that $\textnormal{dad}_{\mathcal{G}}=f$. For any $R>0$, for the open precompact subset $K={\{(\gamma x,\gamma,x)\in\mathcal{G}\mid\ell(\gamma x,\gamma,x)<R\}}$ of $\mathcal{G}$. Also, $\ell_{\Gamma}(\gamma)=\ell(\gamma x,\gamma,x)<R$, thus $K$ has the form
$\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in E\}$
for a finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$. Then there is an open cover $U_{0},\cdots,U_{f(R)}$ of $\mathcal{G}^{(0)}=X$ such that the set
$F_{i}:=\begin{Bmatrix}&\Big{|}&\text{ there exist }x\in U_{i}\text{ and }\gamma_{n},\cdots,\gamma_{1}\in E\text{ such that }\\ \gamma\in\Gamma&\Big{|}&\gamma=\gamma_{n}\cdots\gamma_{2}\gamma_{1}\text{ and for all }k\in\{1,\cdots,n\},\qquad\\ &\Big{|}&\gamma_{k}\cdots\gamma_{1}x\in U_{i}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\end{Bmatrix}$
is finite for each $i\in\{0,\cdots,f(R)\}$. Therefore, the subgroupoid $\mathcal{G}_{i}$ generated by the set
$\{(\gamma x,\gamma,x)\in K\mid x,\gamma x\in U_{i}\}$
is contained in $\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in F_{i}\}.$ Since $F_{i}$ is finite and $X$ is compact, the set $\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in F_{i}\}$ is precompact, and thus $\mathcal{G}_{i}$ is precompact. Hence, $\textnormal{dad}_{\mathcal{G}}(R)\leq f(R)$.
On the other hand, for any $R>0$, for the finite subset $E=\{\gamma\in\Gamma\mid\ell_{\Gamma}(\gamma)<R\}$ of $\Gamma$, there is an open precompact subset $K=\{(\gamma x,\gamma,x)\in\mathcal{G}\mid x\in X,\gamma\in E\}$ of $\mathcal{G}$. Since $\mathcal{G}$ has the dynamic asymptotic dimension function $\textnormal{dad}_{\mathcal{G}}$, there is an open cover $\{U_{0},\cdots,U_{\textnormal{dad}_{\mathcal{G}}(R)}\}$ of $X$ such that for each $i\in\{0,\cdots,\textnormal{dad}_{\mathcal{G}}(R)\}$, the set
$\{(\gamma x,\gamma,x)\in K\mid x,\gamma x\in U_{i}\}$
is contained in a precompact subgroupoid of $\mathcal{G}$. Then the subgroupoid $\mathcal{G}_{i}$ generated by the set
$\{(\gamma x,\gamma,x)\in\mathcal{G}\mid\gamma\in E,x,\gamma x\in U_{i}\},$
is precompact. For each $i$, let
$F_{i}:=\begin{Bmatrix}&\Big{|}&\text{ there exist }x\in U_{i}\text{ and }\gamma_{n},\cdots,\gamma_{1}\in E\text{ such that }\\ \gamma\in\Gamma&\Big{|}&\gamma=\gamma_{n}\cdots\gamma_{2}\gamma_{1}\text{ and for all }k\in\{1,\cdots,n\},\qquad\\ &\Big{|}&\gamma_{k}\cdots\gamma_{1}x\in U_{i}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\end{Bmatrix}.$
Therefore, the subset
$\{(\gamma x,\gamma,x)\in\mathcal{G}\mid\gamma\in F_{i},x\in U_{i}\}$
is contained in $\mathcal{G}_{i}$, and is thus precompact, which implies that $F_{i}$ is finite. Clearly, $f(R)\leq\textnormal{dad}_{\mathcal{G}}(R)$, which forces $f(R)=\textnormal{dad}_{\mathcal{G}}(R)$. Hence, $\textnormal{dad}_{\mathcal{G}}=f$.
The converse can be completed by using the method analogous to that used above. ∎
4.2. Example: coarse spaces with asymptotic dimension growth
In this subsection, we will discuss the connection between our notion and coarse geometry. More specifically, for a discrete metric space of bounded geometry $X$, we will show that the asymptotic dimension growth for $X$ is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid $G(X)$.
Definition 4.10 ([STY02]).
Let $(X,d)$ be a discrete metric space of bounded geometry. The coarse groupoid $G(X)$ associated to $(X,d)$ is defined as follows:
  1. (1)
    for any $R>0$, we define $E_{R}:=\{(x,y)\in X\times X\mid d(x,y)\leq R\}$;
  2. (2)
    as a topological space, we have $G(X):=\bigcup\limits_{R\geq 0}\overline{E_{R}}$ in $\beta(X\times X)$, the Stone-Čech compactification of $X\times X$;
  3. (3)
    equip $G(X)$ with the weak topology coming from this union, this is a locally compact topology for which each $\overline{E_{R}}$ is compact open subset of $G(X)$;
  4. (4)
    $G(X)^{(0)}=\overline{E_{0}}\cong\beta X$;
  5. (5)
    $r(x,y)=x,s(x,y)=y$, for all $(x,y)\in G(X)$;
  6. (6)
    $E_{R}\circ E_{S}\subseteq E_{R+S}$, for all $R,S\geq 0$.
Remark 4.11.
The coarse groupoid $G(X)$ is $\sigma$-compact, locally compact Hausdorff, principal and étale. Recall from [MW20] that there is a canonical length function $\ell$ on $G(X)$ defined by extending the metric $d:X\times X\rightarrow[0,\infty]$ to $\beta(X\times X)$ and observing that it takes finite values on $G(X)$. The canonical length function $\ell$ on $G(X)$ has the following properties:
  1. (1)
    $\ell$ is continuous since it comes from the metric $d$.
  2. (2)
    $\ell$ is coarse. Since $X\times X$ is dense in $\beta(X\times X)$, for any $R>0$, the open set $\ell^{-1}([0,R))$ is contained in $\overline{\ell^{-1}([0,R))\cap X\times X}$, which is contained in the compact set $\overline{E_{R}}$. Hence, $\ell^{-1}([0,R))$ is precompact, and $\ell$ is proper. Observe that for any precompact subset $K$ of $G(X)$, $K$ must contained in a controlled set $E_{R}$ for some $R$ by the definition of topology of $G(X)$. Thus $\ell$ is controlled.
  3. (3)
    $\ell^{-1}([0,R])=\overline{E_{R}}$ and $\ell^{-1}(\{0\})=\overline{E}_{0}\cong\beta X$.
In the case of coarse groupoids, the following theorem bridges the asymptotic dimension growth with the dynamic asymptotic dimension growth.
Theorem 4.12.
Let $X$ be a discrete metric space of bounded geometry, and let $G(X)$ be the associated coarse groupoid. Let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a non-decreasing function. Then $\textnormal{ad}_{X}\approx f$ if and only if $\textnormal{dad}_{G(X)}\approx f$.
Proof.
Assume that $G(X)$ has dynamic asymptotic dimension function $\textnormal{dad}_{G(X)}$, and $X$ has dynamic asymptotic dimension function $\widecheck{\textnormal{ad}}_{X}$ as defined in Lemma 2.15. Below, we show that $\widecheck{\textnormal{ad}}_{X}\approx\textnormal{dad}_{G(X)}$. First, we need to verify that $\widecheck{\textnormal{ad}}_{X}\preceq\textnormal{dad}_{G(X)}$. For any $R>0$, for the open precompact subset $K=\{z\in G(X)\mid\ell(z)<R\}$ of $G(X)$, then there exists a controlled set $E_{R-\frac{1}{n}}$ for some positive integer $n$ such that $\overline{E_{R-\frac{1}{n}}}\subseteq K$. Observe that $\overline{E_{R-\frac{1}{n}}}$ is a compact open subset of $G(X)$ containing the unit space $G(X)^{(0)}=\beta X$. Let $U_{0},\cdots,U_{\textnormal{dad}_{G(X)}(R)}$ be an open cover of $G(X)^{(0)}=\beta X$ such that for each $i\in\{0,\cdots,\textnormal{dad}_{G(X)}(R)\}$, the set
$\{g\in K\mid s(g),r(g)\in U_{i}\}$
is contained in a precompact subgroupoid of $G(X)$. Therefore, the subgroupoid $\mathcal{G}_{i}$ generated by the set
$\{g\in\overline{E_{R-\frac{1}{n}}}\mid s(g),r(g)\in U_{i}\}$
is precompact. By the definition of the topology on $G(X)$, there exists a controlled set $F\subseteq X\times X$ such that each $\mathcal{G}_{i}$ is contained in $\overline{F}\subseteq G(X)$. Fix $i$, and let $\sim$ denote the equivalence relation on $U_{i}$ induced by $\mathcal{G}_{i}$, so $x\sim y$ if there exists $g\in\mathcal{G}_{i}$ such that $s(g)=x,r(g)=y$. We denote by $\{U^{j}_{i}\mid j\in J_{i}\}$ the equivalence classes for this relation. Define
$\mathcal{U}_{i}:=\{U^{j}_{i}\cap X\mid j\in J_{i}\}.$
Let $\mathcal{U}=\mathcal{U}_{0}\sqcup\cdots\sqcup\mathcal{U}_{\textnormal{dad}_{G(X)}(R)}$, and it is a cover of $X$ since $\{U_{0},\cdots,U_{\textnormal{dad}_{G(X)}(R)}\}$ is a cover of $\beta X$.
We next claim that the cover $\mathcal{U}$ is $F$-bounded and each $\mathcal{U}_{i}$ is $E_{R-\frac{1}{n}}$-separated. Indeed, the equivalence relation on each $U^{j}_{i}\cap X$ induced by $\mathcal{G}_{i}$ is entirely contained in $F$ as $\mathcal{G}_{i}\subseteq\overline{F}$, so $\mathcal{U}$ is $F$-bounded. For each $\mathcal{U}_{i}$, if $(U^{j}_{i}\times U^{k}_{i})\cap E_{R-\frac{1}{n}}\neq\varnothing$ for $j\neq k$, then there exists $(x,y)\in(U^{j}_{i}\times U^{k}_{i})\cap E_{R-\frac{1}{n}}$. Observe that $s(x,y)=y,r(x,y)=x$, so $(x,y)\in\mathcal{G}_{i}$ and $x\sim y$. However, when $j\neq k$, $U^{j}_{i}\cap U^{k}_{i}=\varnothing$, contradiction. Thus, for $j\neq k$, $(U^{j}_{i}\times U^{k}_{i})\cap E_{R}=\varnothing$, i.e. $\mathcal{U}_{i}$ is $E_{R-\frac{1}{n}}$-separated. Hence, $\widecheck{\textnormal{ad}}_{X}(R-\frac{1}{n})\leq\textnormal{dad}_{G(X)}(R)$, which implies that $\widecheck{\textnormal{ad}}_{X}\preceq\textnormal{dad}_{G(X)}$.
The next step is to show that $\textnormal{dad}_{G(X)}\preceq\widecheck{\textnormal{ad}}_{X}$. For any $R>0$, for the open precompact subset $K=\{z\in G(X)\mid\ell(z)<R\}$ of $G(X)$. The definition of the topology on $G(X)$ implies that $K\subseteq\overline{E_{R}}$ for the controlled set $E_{R}$. Since $X$ has asymptotic dimension function $\widecheck{\textnormal{ad}}_{X}$, there exist a controlled set $F$ and a cover $\mathcal{U}$ of $X$ such that $\mathcal{U}$ is $F$-bounded and such that
$\mathcal{U}=\mathcal{U}_{0}\sqcup\cdots\sqcup\mathcal{U}_{\widecheck{\textnormal{ad}}_{X}(R)}$
such that each $\mathcal{U}_{i}$ is $E_{R}$-separated. For each $i$, set
$U_{i}:=\overline{\bigsqcup\limits_{U\in\mathcal{U}_{i}}U},$
which is a compact open subset of $\beta X$. Since $\mathcal{U}$ is a cover of $X$, $\{U_{0},\cdots,U_{\widecheck{\textnormal{ad}}_{X}(R)}\}$ is a cover of $\beta X$. Furthermore, for each $i$,
$\mathcal{G}_{i}:=\bigsqcup\limits_{U\in\mathcal{U}_{i}}U\times U$
is a subgroupoid of the pair groupoid that is contained in $F$, and by continuity of the groupoid operations, the subgroupoid of $G(X)$ generated by
$\{g\in E_{R}\mid r(g),s(g)\in U_{i}\}$
is contained in $\overline{\mathcal{G}_{i}}$, a compact subgroupoid of $G(X)$ contained in the compact subset $\overline{F}$. Therefore, $\textnormal{dad}_{G(X)}(R)\leq\widecheck{\textnormal{ad}}_{X}(R)$, which yields that $\textnormal{dad}_{G(X)}\preceq\widecheck{\textnormal{ad}}_{X}$. Hence, we obtain $\textnormal{dad}_{G(X)}\approx\widecheck{\textnormal{ad}}_{X}$ as desired.
Finally, by Lemma 2.15, we see that $\textnormal{ad}_{X}\approx\widecheck{\textnormal{ad}}_{X}$, which implies that $\textnormal{ad}_{X}\approx\textnormal{dad}_{G(X)}$. Therefore, $\textnormal{ad}_{X}\approx f$ if and only if $\textnormal{dad}_{G(X)}\approx f$. ∎
Remark 4.13.
Based on the second step of the above proof, we see that if $\textnormal{ad}_{X}\preceq f$, then $\textnormal{dad}_{G(X)}\preceq f$ in the following slightly stronger form: For any $R>0$, for the open precompact subset $K=\{z\in G(X)\mid\ell(z)<R\}$ of $G(X)$, there exists an open cover $\{U_{0},\cdots,U_{\widecheck{\textnormal{ad}}_{X}(R)}\}$ of $G(X)^{(0)}$ such that the subset $\{g\in K\mid s(g),r(g)\in U_{i}\}$ is contained in a compact open subgroupoid of $G(X)$.
By a result of Ozawa and Oppenheim, the subexponential asymptotic dimension growth implies property A.
Theorem 4.14.
[Oza12, Opp14] Let $X$ be a discrete metric space of bounded geometry. If $X$ has subexponential asymptotic dimension growth, then $X$ has property A.
Combining Theorem 4.12 and Theorem 4.14, we have the following result:
Corollary 4.15.
Let $X$ be a discrete metric space of bounded geometry. If the associated coarse groupoid $G(X)$ has subexponential dynamic asymptotic dimension growth, then $G(X)$ is amenable.
The following example follows from Theorem 4.12 and Corollary 4.15.
Example 4.16.
Let $X$ be a geodesic uniformly locally finite coarse median space with finite rank and at most exponential volume growth, then its corresponding coarse groupoid $G(X)$ has subexponential dynamic asymptotic dimension growth, and it is amenable.
4.3. Equivalence with asymptotic dimension growth
In this subsection, we will build the equivalence between the asymptotic dimension growth for countable discrete groups and the dynamic asymptotic dimension growth for group actions by resorting to the BLR condition. Using this result, we can obtain that countable discrete groups with subexponential asymptotic dimension growth all admit free, amenable, minimal actions on the Cantor set which have at most subexponential dynamic asymptotic dimension growth.
Theorem 4.17.
Let $\Gamma$ be a countable discrete group, and let $h:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function with subexponential growth. Then the following are equivalent:
  1. (1)
    $\Gamma$ admits an action $\Gamma\curvearrowright X$ on a compact space $X$ that is $\textnormal{ad}_{X}$-BLR with respect to the family of finite subgroups and $\textnormal{ad}_{X}\preceq h$;
  2. (2)
    $\Gamma$ admits an action $\Gamma\curvearrowright X$ on a compact space $X$ with dynamic asymptotic dimension function $\textnormal{dad}_{\Gamma\curvearrowright X}\preceq h$.
  3. (3)
    The canonical action of $\Gamma$ on $\beta\Gamma$ has dynamic asymptotic dimension function $\textnormal{dad}_{\Gamma\curvearrowright\beta\Gamma}\preceq h$.
  4. (4)
    $\Gamma$ equipped with the canonical left-invariant coarse structure for which the controlled sets are
    =$\mathcal{E}=\{E\subseteq\Gamma\times\Gamma\mid\{s^{-1}t\in\Gamma\mid(s,t)\in E\}\text{ is finite}\},$
    has asymptotic dimension function $\textnormal{ad}_{X}\preceq h$.
  5. (5)
    The canonical action of $\Gamma$ on $\beta\Gamma$ is $\textnormal{ad}_{X}$-BLR with respect to the family of finite subgroups and $\textnormal{ad}_{X}\preceq h$.
Proof.
The fact that $(1)$ implies $(2)$ follows from Theorem 3.14.
$(2)\Rightarrow(3)$: assume that $\Gamma\curvearrowright X$ is an action on a compact space $X$ with dynamic asymptotic dimension function $\textnormal{dad}_{\Gamma\curvearrowright X}$ and $\textnormal{dad}_{\Gamma\curvearrowright X}\preceq h$. For any $x\in X$, let $f:\Gamma\rightarrow X,\gamma\mapsto\gamma\cdot x$ be any orbit map, and it is $\Gamma$-equivariant. Since $\Gamma$ acts by homeomorphisms on $X$, $f$ is continuous. By universal property of $\beta\Gamma$, $f$ extends uniquely to a continuous map $\phi:\beta\Gamma\rightarrow X$ such that $\phi\circ i=f$, where the inclusion map $i:\Gamma\hookrightarrow\beta\Gamma$ is continuous. Moreover, $\phi$ is $\Gamma$-equivariant as $f$ is $\Gamma$-equivariant and $\Gamma$ is dense in $\beta\Gamma$. By definition, for any $R>0$, for the finite subset $E=\{\gamma\in\Gamma\mid\ell(\gamma)<R\}$ of $\Gamma$, there exists an open cover $\{U_{0},\cdots,U_{\textnormal{dad}_{\Gamma\curvearrowright X}(R)}\}$ of $X$ such that for each $i\in\{0,\cdots,\textnormal{dad}_{\Gamma\curvearrowright X}(R)\}$, the set
[…]
is finite. Note that $\phi(\beta\Gamma)$ is a compact subset of $X$ as $\phi$ is continuous and $\beta\Gamma$ is compact, thus $\{U_{0},\cdots,U_{\textnormal{dad}_{\Gamma\curvearrowright X}(R)}\}$ is also an open cover of $\phi(\beta\Gamma)$. Moreover, $\{\phi^{-1}(U_{0}),\cdots,\phi^{-1}(U_{\textnormal{dad}_{\Gamma\curvearrowright X}(R)})\}$ is an open cover of $\beta\Gamma$ since $\beta\Gamma\subseteq\phi^{-1}(\phi(\beta\Gamma))$ and $\phi$ is continuous. Furthermore, if $y\in\phi^{-1}(U_{i})$ for some $i$ and $\gamma=\gamma_{n}\cdots\gamma_{2}\gamma_{1}$ with $\gamma_{k}\in E$ and $\gamma_{k}\cdots\gamma_{1}y\in\phi^{-1}(U_{i})$ for each $k\in\{1,\cdots,n\}$, then $\phi(y)\in U_{i}$ and $\gamma_{k}\cdots\gamma_{1}\phi(y)=\phi(\gamma_{k}\cdots\gamma_{1}y)\in U_{i}$ as $\phi$ is $\Gamma$-equivariant.
Hence, for each $i\in\{0,\cdots,\textnormal{dad}_{\Gamma\curvearrowright X}(R)\}$, the set
[…]
is contained in the set in line 4.2, and is thus finite. This shows that $\textnormal{dad}_{\Gamma\curvearrowright\beta\Gamma}(R)\leq\textnormal{dad}_{\Gamma\curvearrowright X}(R)$, and thus $\textnormal{dad}_{\Gamma\curvearrowright\beta\Gamma}\leq\textnormal{dad}_{\Gamma\curvearrowright X}\preceq h$.
$(3)\Rightarrow(4)$: the transformation groupoid $\beta\Gamma\rtimes\Gamma$ has dynamic asymptotic dimension function $\textnormal{dad}_{\beta\Gamma\rtimes\Gamma}\preceq h$ by Lemma 4.9. On the other hand, $\beta\Gamma\rtimes\Gamma=G(\Gamma)$ by [Proposition 3.4, [STY02]], thus $G(\Gamma)$ has asymptotic dimension function $\textnormal{dad}_{G(X)}\preceq h$. By Theorem 4.12, $\Gamma$ has asymptotic dimension function $\textnormal{ad}_{X}\preceq h$.
$(4)\Rightarrow(5)$: Combining Theorem 4.14 and the assumption that $\Gamma$ has at most subexponential asymptotic dimension growth, we have that $\Gamma$ has property A. Then the canonical action of $\Gamma$ on $\beta\Gamma$ is amenable by [Theorem 3.3, [HR00]]. Further, given $R>0$, the finite subset $E=\{\gamma\in\Gamma\mid\ell(\gamma)<R\}$ and $\varepsilon>0$, there exists a continuous $f:\beta\Gamma\rightarrow\textnormal{Prob}_{\textnormal{ad}_{X}(R)}(\Gamma)$ such that
$\sup\limits_{\gamma^{\prime}\in\beta\Gamma}\|f(\gamma\gamma^{\prime})-\gamma f(\gamma^{\prime})\|_{1}<\varepsilon,\forall\gamma\in E,$
which implies that $f$ is a $(E,\varepsilon)$-equivariant map. Therefore, the action of $\Gamma$ on $\beta\Gamma$ is $\textnormal{ad}_{X}$-BLR with respect to the family of finite subgroups and $\textnormal{ad}_{X}\preceq h$.
$(5)\Rightarrow(1)$: trivial. ∎
Combining Theorem 4.17 with the results of Rørdam and Sierakowski in [Section 6, [RS12]], we obtain the following theorem.
Theorem 4.18.
Let $\Gamma$ be a countable infinite discrete group with subexponential asymptotic dimension growth $h$. Then $\Gamma$ admits a free, minimal action on the Cantor set which has dynamic asymptotic dimension growth at most $h$.
Proof.
Step 1. In the first step, we will construct countable sets $\{V_{n}\}_{n\in\mathbb{N}}$. From the fact that $(4)$ implies $(5)$ of Theorem 4.17, we see that the action of $\Gamma$ on $\beta\Gamma$ is amenable and $\textnormal{ad}_{X}$-BLR for the family $\mathcal{F}$ of finite subgroups of $\Gamma$ and $\textnormal{ad}_{X}\approx h$. For each $n\in\mathbb{N}$, let $E_{n}=\{\gamma\in\Gamma\mid\ell(\gamma)<n\}$, which is a finite subset of $\Gamma$. By the properties of the length function $\ell$, there is a nested sequence
$E_{1}\subseteq E_{2}\subseteq E_{3}\subseteq\cdots$
such that $\Gamma=\bigcup\limits_{n}E_{n}$. Then for each $n\in\mathbb{N}$, there exist a $(\Gamma,\mathcal{F})$-complex $C_{n}$ of dimension at most $\textnormal{ad}_{X}(n)$ and a continuous $(E_{n},\frac{1}{n})$-equivariant map $f_{n}:\beta\Gamma\rightarrow C_{n}$. Since $\beta\Gamma$ is compact, we assume by Lemma 3.9 that $f_{n}(\beta\Gamma)\subset\textnormal{Prob}(S_{n})\cap C_{n}$, where $S_{n}$ is a finite subset of $C_{n}$ with cardinality $|S_{n}|\leq\textnormal{ad}_{X}(n)+1$. Further, we replace $C_{n}$ by $C_{n}\cap\textnormal{Prob}(\Gamma\cdot S_{n})$, we may assume that there are only finitely many $\Gamma$-orbits of $S_{n}$, which implies that the vertex set $V_{n}$ of $C_{n}$ is countable.
Step 2. In the next step, we show that $\Gamma$ admits a free and $\textnormal{ad}_{X}$-BLR action on a totally disconnected, metrizable compact space $\hat{A}$. Let $A$ be a unital $\Gamma$-invariant $C^{*}$-subalgebra of $\ell^{\infty}(\Gamma)$ , and let $\hat{A}$ denote the spectrum of $A$. By Gelfand duality, $\hat{A}$ identifies with a quotient space of $\beta\Gamma$ such that the quotient map $q:\beta\Gamma\rightarrow\hat{A}$ is equivariant.
For each $n$, write
$f_{n}(x)=\sum\limits_{v\in V_{n}}t_{v}(x)v,$
where each $t_{v}:\beta\Gamma\rightarrow[0,1]$ is a continuous map. Define $\Omega_{1}:=\{t_{v}\mid v\in V_{n},n\in\mathbb{N}\}$, which is a countable subset of $\ell^{\infty}(\Gamma)$ as $\ell^{\infty}(\Gamma)$ identifies with $C(\beta\Gamma)$. If $A$ contains $\Omega_{1}$, then the action of $\Gamma$ on $\hat{A}$ is $\textnormal{ad}_{X}$-BLR. Indeed, the maps $t_{v}:\beta\Gamma\rightarrow[0,1]$ descend to the maps $\hat{A}\rightarrow[0,1]$, whence the maps $f_{n}:\beta\Gamma\rightarrow C_{n}$ descend to the maps $f_{n}:\hat{A}\rightarrow C_{n}$ by [Lemma 3.5, [Hig00]]. On the other hand, [Lemma 6.4, [RS12]] shows that there exists a countable subset $\Omega_{2}$ of $\ell^{\infty}(\Gamma)$ such that if $A$ contains $\Omega_{2}$, then the action of $\Gamma$ on $\hat{A}$ is free.
Let $\Omega=\Omega_{1}\cup\Omega_{2}$, a countable subset of $\ell^{\infty}(\Gamma)$. According to [Lemma 6.7, [RS12]], there exists a countable $\Gamma$-invariant collection of projections $\mathcal{P}$ such that the $C^{*}$-algebra $C^{*}(\mathcal{P})$ generated by $\mathcal{P}$ contains $\Omega$. Let $A=C^{*}(\mathcal{P},1_{\ell^{\infty}(\Gamma)})$, then $\hat{A}$ is a $\Gamma$-space, and the action of $\Gamma$ is free and $\textnormal{ad}_{X}$-BLR. Since $A$ is unital and generated by countably many projections, the space $\hat{A}$ is compact, metrizible and totally disconnected.
Step 3. In this step, we prove that $\Gamma$ admits a free, minimal, $\textnormal{ad}_{X}$-BLR action on a totally disconnected, metrizable compact space $X$. Let $Y:=\hat{A}$, then $\Gamma$ admits a free, $\textnormal{ad}_{X}$-BLR action on a totally disconnected, metrizable compact space $Y$. Observe that if $Z$ is any non-empty closed $\Gamma$-invariant subset of $Y$, then $Z$ and the induced $\Gamma$ action on $Z$ have the above properties. Indeed, there is a nested sequence of closed non-empty $\Gamma$-invariant subsets of $Y$:
$\cdots\subseteq Y_{3}\subseteq Y_{2}\subseteq Y_{1}=Z\subseteq Y.$
Compactness of $Y$ implies that any descending chain for this order has a non-empty intersection. Zorn’s lemma yields that there is a minimal element $X$. Thus the induced $\Gamma$ action on $X$ is minimal. Hence, $\Gamma$ admits a free, minimal, $\textnormal{ad}_{X}$-BLR action on a totally disconnected, metrizable compact space $X$. By Theorem 3.14, this action has dynamic asymptotic dimension function $\textnormal{dad}_{\Gamma\curvearrowright X}\leq\textnormal{ad}_{X}\preceq h$, and thus has dynamic asymptotic dimension growth at most $h$.
Step 4. In the final step, we finish the proof by showing that this $X$ is a copy of the Cantor set, and it thus suffices to show that it has no isolated points. Observe that if $x\in X$ is isolated, then the orbit of $\Gamma\cdot x$ is open, and is dense in $X$ by minimality, thus $X\backslash\Gamma\cdot x=\varnothing$. Since the action is free and $\Gamma$ is infinite, the open cover $\{\{\gamma x\}\mid\gamma\in\Gamma\}$ of $X=\Gamma\cdot x$ has no finite subcover which contradicts compactness. It follows that there is no isolated point. ∎
5. Partitions of unity
In this section, we develop a critical tool that is not only a partition of unity for slow dynamic asymptotic dimension growth, but also provides good properties and estimates. Therefore, this tool plays a key role in proving the amenability of groupoids in the next section. Throughout this section, $\mathcal{G}$ is an étale groupoid, and $\ell$ is a coarse continuous length function on $\mathcal{G}$.
We are now ready to formulate our generalization of [[GWY17], Proposition 7.1].
Proposition 5.1.
Let $\mathcal{G}$ be an étale groupoid with compact unit space, and with dynamic asymptotic dimension function $f\preceq x^{\alpha}(0<\alpha<1)$. Then for any $R>0$, for the open precompact subset $K={\{z\in\mathcal{G}\mid\ell(z)<R\}}$ of $\mathcal{G}$ and any $\varepsilon>0$, there exist a constant $c$ depending on parameters $\alpha,R,\varepsilon$, a positive integer $p$ depending only on $\alpha$, and an open cover $\{U_{0},\cdots,U_{f([(R+1)^{c}]R+1)}\}$ of $\mathcal{G}^{(0)}$ with the following properties:
  1. (1)
    For each $i$, the set
    $\{g\in K\mid s(g),r(g)\in U_{i}\}$
    is contained in an ( open and ) precompact subgroupoid of $\mathcal{G}$.
  2. (2)
    For all $x\in\mathcal{G}^{(0)}$, the ‘partial orbit’ $s(K\cap r^{-1}(x))$ is completely contained in some $U_{i}$.
  3. (3)
    There exists a collection of continuous functions $\{\phi_{i}:\mathcal{G}^{(0)}\rightarrow[0,1]\}_{i=0}^{f([(R+1)^{c}]R+1)}$ such that supp ($\textnormal{supp }(\phi_{i})\subseteq U_{i}$ for each $i$, and $\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\left(\phi_{i}(x)\right)^{p}=1$ for all $x\in\mathcal{G}^{(0)}$, and such that for any $g\in K$,
    $p\cdot\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\left|\phi_{i}(r(g))-\phi_{i}(s(g))\right|<\varepsilon.$
To clearly understand the above proposition, we look at the case of group actions on compact spaces.
Corollary 5.2.
Let $\Gamma\curvearrowright X$ be an action with $X$ compact, and with dynamic asymptotic dimension function $f\preceq x^{\alpha}(0<\alpha<1)$. Then for any $R>0$, for the finite subset $E=\{\gamma\in\Gamma\mid\ell(\gamma)<R\}$ of $\Gamma$ and any $\varepsilon>0$, there exist a constant $c$ depending on parameters $\alpha,R,\varepsilon$, a positive integer $p$ depending only on $\alpha$, and an open cover $\{U_{0},\cdots,U_{f([(R+1)^{c}]R+1)}\}$ of $X$ with the following properties:
  1. (1)
    For each $i$, the set
[…]
    is finite.
  2. (2)
    For all $x\in X$, the collection $E\cdot x:=\{\gamma^{-1}x\mid\gamma\in E\}$ is completely contained in some $U_{i}$.
  3. (3)
    There exists a collection of continuous functions $\{\phi_{i}:\mathcal{G}^{(0)}\rightarrow[0,1]\}_{i=0}^{f([(R+1)^{c}]R+1)}$ such that supp ($\textnormal{supp }(\phi_{i})\subseteq U_{i}$ for each $i$, and $\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\left(\phi_{i}(x)\right)^{p}=1$ for all $x\in X$, and such that for any $\gamma\in E$,
    $p\cdot\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\left|\phi_{i}(\gamma x)-\phi_{i}(x)\right|<\varepsilon.$
To obtain the main proposition of this section, we need several lemmas. The first step is to show the following lemma, which plays an important auxiliary role in the proof of subsequent lemmas and theorems.
Lemma 5.3.
Let $\mathcal{G}$ be an étale groupoid, and let $K$ be a symmetric subset of $\mathcal{G}$ containing $\mathcal{G}^{(0)}$. Let $\mathcal{U}=\{U_{j}\}_{j\in J}$ be an open cover of $\mathcal{G}^{(0)}$. Then for any $g\in K$, $s(g)\in U_{j}$ if and only if $r(g)\in U_{j}$ for some $j\in J$.
Proof.
For any $x\in U_{j}\subseteq\mathcal{G}^{(0)}\subseteq K$, $s(x)=r(x)=x$, thus $U_{j}\subseteq K\cap r^{-1}(U_{j})$, which implies that $U_{j}=s(U_{j})\subseteq s(K\cap r^{-1}(U_{j}))$. If $g\in K$ and $s(g)\in U_{j}$ for some $j$, then $r(g)$ must be in $U_{j}$. Otherwise, $g\notin K\cap r^{-1}(U_{j})$, and so $s(g)\notin s(K\cap r^{-1}(U_{j}))\supseteq U_{j}$, hence $s(g)\notin U_{j}$, contradiction. Conversely, assume that $g\in K$ and $r(g)\in U_{j}$ for some $j$. Due to symmetry, $g^{-1}\in K$, $s(g^{-1})=r(g)\in U_{j}$. From the previous steps, we see that $s(g)=r(g^{-1})\in U_{j}$. ∎
Applying the above conclusion, we have the following lemma, which is an adaption of [[GWY17], Lemma 7.3].
Lemma 5.4.
Let $\mathcal{G}$ be an étale groupoid with compact unit space, and with dynamic asymptotic dimension growth $f$. Then for any $R>0$, for the open precompact subset $K={\{z\in\mathcal{G}\mid\ell(z)<R\}}$ of $\mathcal{G}$, there is a cover $\{U_{0},\cdots,U_{f(R)}\}$ of $\mathcal{G}^{(0)}$ by open precompact subsets of $\mathcal{G}^{(0)}$ such that for each $i$, the set
$\{g\in K\mid s(g),r(g)\in U_{i}\}$
generates an open precompact subgroupoid of $\mathcal{G}$, and moreover so that for each $x\in\mathcal{G}^{(0)}$, the set $s(K\cap r^{-1}(x))$ is completely contained in some $U_{i}$.
Proof.
Since $\mathcal{G}$ has dynamic asymptotic dimension growth $f$ and $\mathcal{G}^{(0)}$ is compact, there is an open cover $\{U_{0},\cdots,U_{f(R)}\}$ of $\mathcal{G}^{(0)}$ by precompact subsets of $\mathcal{G}^{(0)}$ such that for each $i$, the set
$\{g\in K\mid s(g),r(g)\in U_{i}\}$
is contained in a precompact subgroupoid of $\mathcal{G}$, and hence generates an open precompact subgroupoid of $\mathcal{G}$. Observe that $K$ is symmetric as $\ell(x)=\ell(x^{-1})$ for all $x\in\mathcal{G}$. Moreover, for all $x\in\mathcal{G}^{(0)}\subseteq K$, we know that $r(x)=x$, so $x\in K\cap r^{-1}(x)$, thus $K\cap r^{-1}(x)\neq\varnothing$. For any $g\in K\cap r^{-1}(x)$, we have $g\in K$ and $r(g)=x\in U_{i}$ for some $i$. By Lemma 5.3, $s(g)\in U_{i}$. Hence, $s(K\cap r^{-1}(x))\subseteq U_{i}$ for some $U_{i}$. ∎
If the unit space of an étale groupoid is compact and Hausdorff, then it is normal, which yields an excellent result as follows.
Lemma 5.5.
Let $\mathcal{G}$ be an étale groupoid with compact unit space, and let $V_{0},\cdots,V_{n}$ be an open cover of $\mathcal{G}^{(0)}$. Then there exist open subsets $\mathcal{U}_{0},\cdots,\mathcal{U}_{n}$ of $\mathcal{G}^{(0)}$ such that $\bigcup\limits_{i=0}^{n}\mathcal{U}_{i}$ covers $\mathcal{G}^{(0)}$ and $\overline{\mathcal{U}_{i}}\subseteq V_{i}$ for each $i$.
Proof.
We divide our proof into four steps. First, we show that for any open subset $V$ of $\mathcal{G}^{(0)}$ and any $x\in V$, there exists an open neighborhood $U_{x}$ of $x$ such that $\overline{U_{x}}\subseteq V$. Note that $\mathcal{G}^{(0)}$ is Hausdorff, thus the singleton set $\{x\}$ is closed. Also, $\mathcal{G}^{(0)}$ is compact, then it is normal. Observe that the boundary point $\partial V$ is closed, and $\{x\}\cap\partial V\subseteq V\cap\partial V=\varnothing$ as $V$ is open. Hence by normality of $\mathcal{G}^{(0)}$, there exist an open neighborhood $U_{x}$ of $x$ and open set $V^{\prime}$ containing $\partial V$ such that
$U_{x}\cap V^{\prime}=\varnothing.$
We may assume that $U_{x}\subseteq V$, otherwise we replace $U_{x}$ by $U_{x}\cap V$. It suffices to show that $\overline{U_{x}}\subseteq V$. If this does not happen, we may find $y\in\partial U_{x}\subseteq\overline{V}$ such that $y\notin V$, which implies $y\in\partial V\subseteq V^{\prime}$. It follows that $V^{\prime}$ is an open neighborhood of $y$, and that $U_{x}\cap V^{\prime}\neq\varnothing$ as $y\in\partial U_{x}$, contradiction. Hence, $\overline{U_{x}}\subseteq V$ as required.
In the second step, we build subsets $V_{0}^{\prime}$ and $\mathcal{U}_{0}$ of $\mathcal{G}^{(0)}$ such that
$V_{0}^{\prime}\subseteq\mathcal{U}_{0}\subseteq\overline{\mathcal{U}_{0}}\subseteq V_{0}.$
If $y\in\partial V_{0}$, namely $y\notin V_{0}$, then $y$ must be in $V_{i}$ for some $i\in\{1,\cdots,n\}$. Thus there exists an open neighborhood $U_{y}^{\prime}$ of $y$ such that $U_{y}^{\prime}\subseteq V_{i}$ as $V_{i}$ is open. Observe that for any $z\in\partial V_{i}$, $z$ and $y$ are distinct points. Since $\mathcal{G}^{(0)}$ is Hausdorff, there exist open neighborhoods $\widetilde{U_{z}}$ of $z$ and $\widetilde{U_{y}}$ of $y$ such that $\widetilde{U_{z}}\cap\widetilde{U_{y}}=\varnothing$. Set $U_{y}=\widetilde{U_{y}}\cap U_{y}^{\prime}$, then $U_{y}\cap\widetilde{U_{z}}=\varnothing$ and $U_{y}\subseteq V_{i}$. Let $V_{0}^{\prime}=\overline{V_{0}}\backslash\bigcup\limits_{y\in\partial V_{0}}U_{y}$, it is closed and hence compact. Note that $V_{0}^{\prime}\subseteq V_{0}$, then by the first step, for any $x\in V_{0}^{\prime}$, there exists an open neighborhood $U_{x}$ of $x$ such that $\overline{U_{x}}\subseteq V_{0}$ and $\bigcup\limits_{x\in V_{0}^{\prime}}U_{x}$ is a cover of $V_{0}^{\prime}$. In view of the compactness of $V_{0}^{\prime}$, we can find finitely many points $x_{0}^{0},\cdots,x_{0}^{m}$ in $V_{0}^{\prime}$ such that $V_{0}^{\prime}\subseteq\bigcup\limits_{k=0}^{m}U_{x_{0}^{k}}$ with $\overline{U_{x_{0}^{k}}}\subseteq V_{0}$. Let $\mathcal{U}_{0}=\bigcup\limits_{k=0}^{m}U_{x_{0}^{k}}$, then $V_{0}^{\prime}\subseteq\mathcal{U}_{0}$ and $\overline{\mathcal{U}_{0}}=\overline{\bigcup\limits_{k=0}^{m}U_{x_{0}^{k}}}=\bigcup\limits_{k=0}^{m}\overline{U_{x_{0}^{k}}}\subseteq V_{0}$.
Another step is to construct subsets $V_{i}^{\prime}$ and $\mathcal{U}_{i}$ of $\mathcal{G}^{(0)}$ such that $U_{y}\subseteq V_{i}^{\prime}$ when $y\in\partial V_{0}\cap V_{i}$ for some $i\in\{1,\cdots,n\}$, and such that $V_{i}^{\prime}\subseteq\mathcal{U}_{i}\subseteq\overline{\mathcal{U}_{i}}\subseteq V_{i}$. From the second step, we see that $U_{y}\subseteq V_{i}$ and $U_{y}\cap\widetilde{U_{z}}=\varnothing$, $\forall z\in\partial V_{i}$. For fixed $i$ and $z\in\partial V_{i}$, $z\notin V_{i}$ implies that there is at least one $l\in\{0,\cdots,i-1,i+1,\cdots,n\}$ such that $z\in V_{l}$. Therefore, there exists an open neighborhood $U_{z}^{\prime}$ of $z$ such that $U_{z}^{\prime}\subseteq V_{l}$ as $V_{l}$ is open. Similar to the second step, for any $h\in\partial V_{l}$, there exist open neighborhoods $\widetilde{U_{h}}$ of $h$ and $U_{z}^{\prime\prime}$ of $z$ such that $\widetilde{U_{h}}\cap U_{z}^{\prime\prime}=\varnothing$. Replacing $U_{z}^{\prime}$ by $U_{z}^{\prime}\cap U_{z}^{\prime\prime}$, we have $\widetilde{U_{h}}\cap U_{z}^{\prime}=\varnothing$ and $U_{z}^{\prime}\subseteq V_{l}$. Set $U_{z}=\widetilde{U_{z}}\cap U_{z}^{\prime}$, then $U_{z}\subseteq V_{l}$, $U_{z}\cap U_{y}=\varnothing$ and $U_{z}\cap\widetilde{U_{h}}=\varnothing$. Let $V_{i}^{\prime}=\overline{V_{i}}\backslash\bigcup\limits_{z\in\partial V_{i}}U_{z}$, by $U_{z}\cap U_{y}=\varnothing$, then we get the desired inclusion $U_{y}\subseteq V_{i}^{\prime}$. Repeating the second step, we may obtain an open subset $\mathcal{U}_{i}$ such that $V_{i}^{\prime}\subseteq\mathcal{U}_{i}\subseteq\overline{\mathcal{U}_{i}}\subseteq V_{i}$.
Finally, we show that there are open subsets $\mathcal{U}_{0},\cdots,\mathcal{U}_{n}$ of $\mathcal{G}^{(0)}$ covering $\mathcal{G}^{(0)}$ such that $\overline{\mathcal{U}_{i}}\subset V_{i}$ for each $i$. By the third step, for any $y\in\partial V_{0}$, there exist $i\in\{1,\cdots,n\}$ and a closed subset $V_{i}^{\prime}$ such that $U_{y}\subseteq V_{i}^{\prime}$. Therefore, $\bigcup\limits_{y\in\partial V_{0}}U_{y}\subseteq\bigcup\limits_{i=1}^{n}V_{i}^{\prime}\subseteq\bigcup\limits_{i=1}^{n}\mathcal{U}_{i}$. It follows that $\overline{V_{0}}=V_{0}^{\prime}\cup\bigcup\limits_{y\in\partial V_{0}}U_{y}\subseteq\mathcal{U}_{0}\cup\bigcup\limits_{i=1}^{n}\mathcal{U}_{i}=\bigcup\limits_{i=0}^{n}\mathcal{U}_{i}$. Repeating the previous steps, we obtain
$\overline{V_{j}}\subseteq\bigcup\limits_{i=0}^{n}\mathcal{U}_{i},\text{for each }j\in\{0,\cdots,n\}.$
Hence, $\mathcal{G}^{(0)}\subseteq\bigcup\limits_{j=0}^{n}\overline{V_{j}}\subseteq\bigcup\limits_{i=0}^{n}\mathcal{U}_{i}$ with $\overline{\mathcal{U}_{i}}\subseteq V_{i}$ for each $i$. ∎
With the help of the preceding two lemmas, we can now acquire the following key lemma, which is an adaption of [[GWY17], Lemma 7.4].
Lemma 5.6.
Let $\mathcal{G}$ be an étale groupoid with compact unit space, and with dynamic asymptotic dimension function $f$. Let $N:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a nondecreasing function. For any $R>0$, for the open precompact subset $K={\{z\in\mathcal{G}\mid\ell(z)<R\}}$ of $\mathcal{G}$, and each $i\in\{0,\cdots,f(N(R)R+1)\}$, there is a nested family
$U^{(0)}_{i}\subseteq U^{(1)}_{i}\subseteq\cdots\subseteq U_{i}^{(N(R))}$
of open precompact subsets of $\mathcal{G}^{(0)}$ with the following properties.
  1. (1)
    The family $\{U_{0}^{(0)},\cdots,U_{f(N(R)R+1)}^{(0)}\}$ covers $\mathcal{G}^{(0)}$;
  2. (2)
    For all $i,n$, $\overline{U^{(n)}_{i}}\subseteq U_{i}^{(n+1)}$;
  3. (3)
    For all $i,n$,
    $s(K\cap r^{-1}(U^{(n)}_{i}))\subseteq U_{i}^{(n+1)};$
  4. (4)
    For all $i$, the set
    $\{g\in K\mid s(g),r(g)\in U_{i}^{(N(R))}\}$
    generates an open precompact subgroupoid of $\mathcal{G}$.
Proof.
We will prove this lemma in two steps. First, we will build a nested sequence of open subsets of $\mathcal{G}^{(0)}$. Given $R>0$ and a nondecreasing function $N:\mathbb{R}^{+}\rightarrow\mathbb{N}$. Observe that $K$ is symmetric and contains $\mathcal{G}^{(0)}$. For each $n\in\mathbb{N}$, we define
$\overline{K}^{n}=\{g_{n}\cdots g_{1}\mid g_{k}\in\overline{K}\text{ and }s(g_{k+1})=r(g_{k}),\forall k\in\{1,\cdots,n-1\}\}.$
Since $\overline{K}$ is compact, we see that $\overline{K}^{n}$ is compact by Lemma 4.8. Also, $\overline{K}^{n}\subseteq\overline{K}^{n+1}$ for each $n\in\mathbb{N}$ as long as we let $g_{n+1}=r(g_{n})\in\overline{K}$. Observe that for any $z\in\overline{K}^{N(R)}$, $\ell(z)\leq N(R)R<N(R)R+1$, thus $\overline{K}^{N(R)}$ is contained in $K^{\prime}=\{z\in\mathcal{G}\mid\ell(z)<N(R)R+1\}$, which is an open precompact symmetric subset of $\mathcal{G}$ containing $\mathcal{G}^{(0)}$. According to Lemma 5.4, there are open precompact subsets $V_{0},\cdots,V_{f(N(R)R+1)}$ of $\mathcal{G}^{(0)}$ that cover $\mathcal{G}^{(0)}$ such that for each $i$, the set $\{g\in K^{\prime}\mid s(g),r(g)\in V_{i}\}$ generates an open precompact subgroupoid of $\mathcal{G}$.
In addition, for each $i\in\{0,\cdots,f(N(R)R+1)\}$ and each $n\in\{0,\cdots,N(R)\}$, we define
$V_{i}^{(n)}:=\{x\in V_{i}\mid s(\overline{K}^{N(R)-n}\cap r^{-1}(x))\subseteq V_{i}\}$
The sequence $\mathcal{G}^{(0)}=\overline{K}^{0}\subseteq\overline{K}\subseteq\cdots\subseteq\overline{K}^{N(R)}$ induces a nested sequence of subsets
$V^{(0)}_{i}\subseteq V_{i}^{(1)}\subseteq\cdots\subseteq V_{i}^{(N(R))}=V_{i}.$
By Lemma 5.4, for any $x\in\mathcal{G}^{(0)}$, $s(K^{\prime}\cap r^{-1}(x))\subseteq V_{i}$ for some $i$, namely, if $g\in K^{\prime}$ with $r(g)=x$, then $s(g)\in V_{i}$. By Lemma 5.3, $x=r(g)\in V_{i}$. Further, $s(\overline{K}^{N(R)}\cap r^{-1}(x))\subseteq s(K^{\prime}\cap r^{-1}(x))\subseteq V_{i}$, then $x\in V_{i}^{(0)}$. Therefore, the family $\{V_{0}^{(0)},\cdots,V^{(0)}_{f(N(R)R+1)}\}$ covers $\mathcal{G}^{(0)}$.
We claim that $V_{i}^{(n)}$ is open. Say for contradiction that $V_{i}^{(n)}$ is not open, then we may find a net $(x_{\lambda})_{\lambda\in\Lambda}$ in $\mathcal{G}^{(0)}$ with $x_{\lambda}\rightarrow x$ for some $x\in V_{i}^{(n)}$ such that for each $\lambda\in\Lambda$, there exists $g_{\lambda}\in\overline{K}^{N(R)-n}\cap r^{-1}(x_{\lambda})$ such that $s(g_{\lambda})\notin V_{i}$. Since $\overline{K}^{N(R)-n}$ is compact, the net $(g_{\lambda})_{\lambda\in\Lambda}$ has a convergent subnet $(g_{\lambda_{j}})_{j\in J}$, thus there exists $g\in\overline{K}^{N(R)-n}$ such that $\lim\limits_{j}g_{\lambda_{j}}=g$. As $r$ is continuous and $r(g_{\lambda})=x_{\lambda}$, $x=\lim\limits_{\lambda}x_{\lambda}=\lim\limits_{j}x_{\lambda_{j}}=\lim\limits_{j}r(g_{\lambda_{j}})=r(g)$, then $g\in\overline{K}^{N(R)-n}\cap r^{-1}(x)$. Because $V_{i}$ is open and $s(g_{\lambda})\notin V_{i}$, $s(g)\notin V_{i}$, which contradicts that $x\in V_{i}^{(n)}$. Hence, $V_{i}^{(n)}$ is open as claimed.
In the second step, we will construct the desired nested sequence $U_{i}^{(n)}$. Since $\mathcal{G}^{(0)}$ is compact, by Lemma 5.5, there exist open subsets $U_{0}^{(0)},\cdots,U_{f(N(R)R+1)}^{(0)}$ of $\mathcal{G}^{(0)}$ covering $\mathcal{G}^{(0)}$ and $\overline{U_{i}^{(0)}}\subseteq V_{i}^{(0)}$ for each $i$. For each $i$,
$\overline{s(K\cap r^{-1}(U_{i}^{(0)}))}\subseteq s(\overline{K}\cap r^{-1}(\overline{U_{i}^{(0)}}))\subseteq s(\overline{K}\cap r^{-1}(V_{i}^{(0)}))\subseteq V_{i}^{(1)}.$
By normality of the compact set $\overline{V_{i}}$, there exists an open set $U_{i}^{(1)}$ such that
$\overline{s(K\cap r^{-1}(U_{i}^{(0)}))}\subseteq U_{i}^{(1)}\subseteq\overline{U_{i}^{(1)}}\subseteq V_{i}^{(1)}.$
We also see that
$\overline{s(K\cap r^{-1}(U_{i}^{(1)}))}\subseteq s(\overline{K}\cap r^{-1}(V_{i}^{(1)}))\subseteq V_{i}^{(2)},$
there exists an open set $U_{i}^{(2)}$ such that
$\overline{s(K\cap r^{-1}(U_{i}^{(1)}))}\subseteq U_{i}^{(2)}\subseteq\overline{U_{i}^{(2)}}\subseteq V_{i}^{(2)}.$
Repeating this process, we get a nested sequence of open subsets
$U_{i}^{(0)}\subseteq U_{i}^{(1)}\subseteq\cdots\subseteq U_{i}^{(N(R))},$
such that for each $i,n$,
$\overline{s(K\cap r^{-1}(U_{i}^{(n)}))}\subseteq U_{i}^{(n+1)}\text{ and }\overline{U_{i}^{(n)}}\subseteq V_{i}^{(n)}.$
Due to $U_{i}^{(n)}\subseteq V_{i}^{(n)}\subseteq V_{i}\subseteq\mathcal{G}^{(0)}$, $U_{i}^{(n)}$ is precompact. Moreover, $\overline{U_{i}^{(n)}}\subseteq\overline{s(K\cap r^{-1}(U_{i}^{(n)}))}\subseteq U_{i}^{(n+1)}$. Note that for all $i$, the set
$\{g\in K\mid s(g),r(g)\in U_{i}^{(N(R))}\}\subseteq\{g\in K^{\prime}\mid s(g),r(g)\in V_{i}\},$
which implies that the set $\{g\in K\mid s(g),r(g)\in U_{i}^{(N(R))}\}$ generates an open precompact subgroupoid of $\mathcal{G}$. ∎
So far, we have built the tools required for Proposition 5.1, and now we can start proving it.
Proof of Proposition 5.1.
Step 1. In the first step, we build a desired open cover of $\mathcal{G}^{(0)}$. Without loss of generality, we may assume that $\mathcal{G}$ has dynamic asymptotic dimension function $f$. Given $R>0$, the open precompact subset $K=\{z\in\mathcal{G}\mid\ell(z)<R\}$ of $\mathcal{G}$, and $\varepsilon>0$. Clearly, $K=K^{-1}$, $s(K)\cup r(K)=\mathcal{G}^{(0)}\subseteq K$. Let $p=[\frac{\alpha}{1-\alpha}]+1$, and let $N(R)=[(R+1)^{c}]$ for some constant $c$ in Lemma 5.6, such that
$\frac{2p\left(f(N(R)R+1)+1\right)+2p\left(f(N(R)R+1)+1\right)^{\frac{1}{p}+1}}{N(R)}<\varepsilon,$ (5.1)
such constant $c$ exists and depends on given parameters $\alpha,R,\varepsilon$. Let $\{U_{i}^{(n)}\mid i\in\{0,\cdots,f(N(R)R+1)\},n\in\{0,\cdots,N(R)\}\}$ have the properties in Lemma 5.6. For each $i$, let $U_{i}=U_{i}^{(N(R))}$. By Lemma 5.6 (i)(ii), $\{U_{0}^{(0)},\cdots,U_{f(N(R)R+1)}^{(0)}\}$ covers $\mathcal{G}^{(0)}$ and $U_{i}^{(0)}\subseteq U_{i}^{(N(R))}=U_{i}$, thus the family $\{U_{0},\cdots,U_{f(N(R)R+1)}\}$ covers $\mathcal{G}^{(0)}$. By Lemma 5.6 (iii), we have that $s(K\cap r^{-1}(U_{i}^{(0)}))\subseteq U_{i}^{(1)}\subseteq U_{i}$ for some $i$, which implies that for any $x\in\mathcal{G}^{(0)}$, $s(K\cap r^{-1}(x))\subseteq U_{i}$ for some $U_{i}$. By Lemma 5.6 (iv), we see that $\{g\in K\mid s(g),r(g)\in U_{i}\}$ generates an open precompact subgroupoid of $\mathcal{G}$.
Step 2. The second step is to construct a partition of unity. By Lemma 5.6 (ii), for each $i\in\{0,\cdots,f(N(R)R+1)\}$, and each $n\in\{1,\cdots,N(R)\}$, $\overline{U_{i}^{(n-1)}}\subseteq U_{i}^{(n)}$. According to Urysohn’s lemma, there exist continuous functions
$\psi_{i}^{(n)}:\mathcal{G}^{(0)}\rightarrow[0,1]$
such that $\psi_{i}^{(n)}(x)=1$, $\forall x\in U_{i}^{(n-1)}$; $\psi^{(n)}_{i}(x)=0$, $\forall x\in\mathcal{G}^{(0)}\backslash U_{i}^{(n)}$.
For $i\in\{0,\cdots,f(N(R)R+1)\}$, put
$\psi_{i}=\frac{1}{N(R)}\sum\limits_{n=1}^{N(R)}\psi_{i}^{(n)},$
then supp ($\textnormal{supp }(\psi_{i})\subseteq U_{i}$ and $\psi_{i}(x)\leq 1,\forall x\in\mathcal{G}^{(0)}$. Define
$\phi_{i}=\frac{\psi_{i}}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}\right)^{\frac{1}{p}}},$
which are continuous functions and supp ($\textnormal{supp }(\phi_{i})\subseteq U_{i}$. By Lemma 5.6 (i), for any $x\in\mathcal{G}^{(0)}$, there exists at least one $j\in\{0,\cdots,f(N(R)R+1)\}$ such that $x\in U_{j}^{(0)}$, thus $\psi_{j}(x)=1$, so $\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(x)\right)^{\frac{1}{p}}\geq 1$, therefore
$10\leq\phi_{i}(x)\leq 1\text{ and }\sum\limits_{i=0}^{f(N(R)R+1)}\left(\phi_{i}(x)\right)^{p}=$ 1 .
Step 3. The final step is to show that for any $g\in K$,
$p\cdot\sum\limits_{i=0}^{f(N(R)R+1)}\left|\phi_{i}(s(g))-\phi_{i}(r(g))\right|<\varepsilon.$
For each $j\in\{0,\cdots,f(N(R)R+1)\}$, set $U_{j}^{(n)}=\mathcal{G}^{(0)}$ for $n\geq N(R)+1$. Define $M=M_{j}:=\min\{n\mid r(g)\in U_{j}^{(n)}\}$. Thus $r(g)\in U_{j}^{(M)}\backslash U_{j}^{(M-1)}$, and
$\psi_{j}^{(n)}(r(g))=\begin{cases}1,\quad&n\in[M+1,N(R)]\\ 0,\quad&n\in[1,M-1]\\ [0,1],\quad&n=M.\end{cases}$
Hence
$\frac{N(R)-(M+1)}{N(R)}\leq\psi_{j}(r(g))\leq\frac{N(R)-M}{N(R)}.$ (5.2)
Since $g\in K$ and $r(g)\in U_{j}^{(M)}$, $g\in K\cap r^{-1}(U_{j}^{(M)})$. By Lemma 5.6 (ii), we see that $s(g)\in s(K\cap r^{-1}(U_{j}^{(M)}))\subseteq U_{j}^{(M+1)}$. Note that $r(g)\notin U_{j}^{(M-1)}\supseteq s(K\cap r^{-1}(U_{j}^{(M-2)}))$, thus $s(g^{-1})=r(g)\notin s(K\cap r^{-1}(U_{j}^{(M-2)}))$. We see that $g^{-1}\in K$ as $K=K^{-1}$, thus $g^{-1}\notin r^{-1}(U_{j}^{(M-2)})$, which implies that $s(g)=r(g^{-1})\notin U_{j}^{(M-2)}$. Therefore $s(g)\in U_{j}^{(M+1)}\backslash U_{j}^{(M-2)}$, and
$\psi_{j}^{(n)}(s(g))=\begin{cases}1,\quad&n\in[M+2,N(R)]\\ 0,\quad&n\in[1,M-2]\\ [0,1],\quad&n=M-1,M,M+1.\end{cases}$
Hence
$\frac{N(R)-(M+2)}{N(R)}\leq\psi_{j}(s(g))\leq\frac{N(R)-(M-1)}{N(R)}.$ (5.3)
Combining equation 5.2 and equation 5.3, we have that
$\left|\psi_{j}(s(g))-\psi_{j}(r(g))\right|\leq\frac{2}{N(R)}.$ (5.4)
By Lemma 5.3, there exists at least one $j\in\{0,\cdots,f(N(R)R+1)\}$ such that $s(g),r(g)$ in $U_{j}^{(0)}$, hence $\psi_{j}(r(g))=\psi_{j}(s(g))=1$, which implies that
$\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(r(g))\geq 1\text{ and }\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(s(g))\geq 1.$ (5.5)
For fixed $i\in\{0,\cdots,f(N(R)R+1)\}$ and $g\in K$,
$\left|\phi_{i}(r(g))-\phi_{i}(s(g))\right|$ $=\left|\frac{\psi_{i}(r(g))}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(r(g))\right)^{\frac{1}{p}}}-\frac{\psi_{i}(s(g))}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(s(g))\right)^{\frac{1}{p}}}\right|$
$\leq\frac{1}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(r(g))\right)^{\frac{1}{p}}}\left|\psi_{i}(r(g))-\psi_{i}(s(g))\right|$
$+\left|\psi_{i}(s(g))\right|\left|\frac{1}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(r(g))\right)^{\frac{1}{p}}}-\frac{1}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(s(g))\right)^{\frac{1}{p}}}\right|.$ (5.6)
Combining equation 5.4 and equation 5.5, we have
$\frac{1}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(r(g))\right)^{\frac{1}{p}}}\left|\psi_{i}(r(g))-\psi_{i}(s(g))\right|\leq\frac{2}{N(R)}.$ (5.7)
On the other hand, using equations 5.4 and 5.5 again, we obtain
$\left|\psi_{i}(s(g))\right|\left|\frac{1}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(r(g))\right)^{\frac{1}{p}}}-\frac{1}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(s(g))\right)^{\frac{1}{p}}}\right|$
$\leq\frac{\left|\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(s(g))\right)^{\frac{1}{p}}-\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(r(g))\right)^{\frac{1}{p}}\right|}{\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(r(g))\right)^{\frac{1}{p}}\left(\sum\limits_{j=0}^{f(N(R)R+1)}\psi_{j}^{p}(s(g))\right)^{\frac{1}{p}}}$
$\leq\left(\sum\limits_{j=0}^{f(N(R)R+1)}\left|\psi_{j}(s(g))-\psi_{j}(r(g))\right|^{p}\right)^{\frac{1}{p}}$
$\leq\frac{2\left(f(N(R)R+1)+1\right)^{\frac{1}{p}}}{N(R)},$ (5.8)
where the third inequality comes from the Minkowski inequality
$\left|\left(\sum\limits_{j=0}^{d}a_{j}^{p}\right)^{\frac{1}{p}}-\left(\sum\limits_{j=0}^{d}b_{j}^{p}\right)^{\frac{1}{p}}\right|\leq\left(\sum\limits_{j=0}^{d}|a_{j}-b_{j}|^{p}\right)^{\frac{1}{p}},\text{ for }a_{j},b_{j}\geq 0.$
According to equations 5, 5.7, and 5, we have
$\left|\phi_{i}(r(g))-\phi_{i}(s(g))\right|$ $\leq\frac{2}{N(R)}+\frac{2\left(f(N(R)R+1)+1\right)^{\frac{1}{p}}}{N(R)}$
$=\frac{2+2\left(f(N(R)R+1)+1\right)^{\frac{1}{p}}}{N(R)}.$
Finally, from the equation 5.1, we obtain
$p\cdot\sum\limits_{i=0}^{f(N(R)R+1)}\left|\phi_{i}(r(g))-\phi_{i}(s(g))\right|$
$\leq\frac{2p\left(f(N(R)R+1)+1\right)+2p\left(f(N(R)R+1)+1\right)^{\frac{1}{p}+1}}{N(R)}<\varepsilon.$
Remark 5.7.
Given $10<\alpha<$ 1, $R>0$ and $\varepsilon>0$. To see the existence of $c$, we consider such example: $f(x)=[x^{\alpha}]$ , $p=[\frac{\alpha}{1-\alpha}]+1$. Through precise calculations, we found that if the constant $c$ satisfies the following inequality
$c>\max\left\{\frac{\alpha(p+1)\textnormal{In}2-p\textnormal{In}\frac{\varepsilon}{4p}}{\left(p-\alpha(p+1)\right)\textnormal{In}(R+1)},\frac{\alpha\textnormal{In}2-\textnormal{In}\frac{\varepsilon}{8p}}{(1-\alpha)\textnormal{In}(R+1)}\right\},$
we have
$\frac{2p\left(f([(R+1)^{c}]R+1)+1\right)+2p\left(f([(R+1)^{c}]R+1)+1\right)^{\frac{1}{p}+1}}{[(R+1)^{c}]}<\varepsilon.$
In fact, the finite dynamic asymptotic dimension also deduces almost invariant partitions of unity without the restriction of the compactness of the unit space, which is a suitable adaption of [Proposition 7.1, [GWY17]]. Hence, we get a stronger result as follows.
Proposition 5.8 ([GWY17]).
Let $\mathcal{G}$ be an étale groupoid with dynamic asymptotic dimension $d$. Then for any open precompact subset $K$ of $\mathcal{G}$ and $\varepsilon>0$, there exists an open cover $\{U_{0},\cdots,U_{d}\}$ of $r(K)\cup s(K)$ with the following properties.
  1. (1)
    For each $i$, the set
    $\{g\in K\mid s(g),r(g)\in U_{i}\}$
    is contained in an ( open and ) precompact subgroupoid of $\mathcal{G}$.
  2. (2)
    For each $x\in r(K)\cup s(K)$, the ‘partial orbit’ $s(r^{-1}(x)\cap K)$ is completely contained in some $U_{i}$.
  3. (3)
    There exists a collection of continuous functions $\{\phi_{i}:\mathcal{G}^{(0)}\rightarrow[0,1]\}_{i=0}^{d}$ such that the support of each $\phi_{i}$ is contained in $U_{i}$, $\sum\limits_{i=0}^{d}\phi_{i}(x)\leq 1$ for all $x\in\mathcal{G}^{(0)}$, and $\sum\limits_{i=0}^{d}\phi_{i}(x)=1$ for all $x\in r(K)\cup s(K)$, and such that for any $g\in K$ and each $i$,
    $|\phi_{i}(r(g))-\phi_{i}(s(g))|<\varepsilon.$
6. Amenability
Our motivation for discussing the relationship between the dynamic asymptotic dimension growth of a groupoid and the amenability is fourfold. First, in the work of [Tu99], Tu showed that the Baum-Connes conjecture holds for foliations whose holonomy groupoid is Hausdorff and amenable, particularly for amenable groupoids. Second, for a countable discrete group $\Gamma$ with property A, Higson showed that the Novikov conjecture holds [Hig00]. If such $\Gamma$ has bounded geometry, then property A is equivalent to the amenability of its associated coarse groupoid $G(\Gamma)$ [STY02]. Therefore, if a countable discrete group with bounded geometry has the amenable coarse groupoid, then the Novikov conjecture holds. Third, according to Corollary 4.15, we see that the subexponential dynamic asymptotic dimension growth of a coarse groupoid implies the amenability. Finally, analogous to the fact that finite dynamical complexity of a groupoid implies the amenability [GWY22], we prove that every locally compact Hausdorff étale groupoid $\mathcal{G}$ with finite dynamic asymptotic dimension is amenable by using the optimized Proposition 5.8. More generally, applying Proposition 5.1 in the previous section, we directly prove that every $\sigma$-compact, locally compact Hausdorff étale groupoid $\mathcal{G}$ with compact unit space having dynamic asymptotic dimension growth at most $x^{\alpha}$ $(0<\alpha<1)$ is amenable.
A reference [AR00] provides a standard description of an amenable groupoid and another reference [BO08] depicts the étale case. The following equivariant definition is a slight variant of [Definition 5.6.13, [BO08]], which is used in the proofs of two theorems of this section.
Definition 6.1 ([GWY22]).
Let $\mathcal{G}$ be a locally compact, Hausdorff, étale groupoid. Say $\mathcal{G}$ is amenable if for all compact subset $K\subseteq\mathcal{G}$ and all $\varepsilon>0$, there exists a continuous, compactly supported function $\mu:\mathcal{G}\rightarrow[0,1]$ such that:
  1. (1)
    for all $x\in\mathcal{G}^{(0)}$, we have $\sum\limits_{g\in\mathcal{G}_{x}}\mu(g)\leq 1$;
  2. (2)
    for all $k\in K$, we have $|1-\sum\limits_{g\in\mathcal{G}_{r(k)}}\mu(g)|<\varepsilon$;
  3. (3)
    for all $k\in K$, we have $\sum\limits_{g\in\mathcal{G}_{r(k)}}|\mu(g)-\mu(gk)|<\varepsilon$.
The following lemma is a vital tool in the proofs of two theorems.
Lemma 6.2 ([GWY22]).
Let $\mathcal{G}$ be an étale groupoid, and let $H$ be an open subgroupoid of $\mathcal{G}$ with compact closure. Then $H$ is amenable.
The following theorem is inspired in part by Guentner,Willett, and Yu’s work [GWY22] on dynamical complexity and controlled operator $K$-theory, and their work [GWY17] on finite dynamic asymptotic dimension. With the help of the nuclear dimension for the reduced groupoid $C^{*}$-algebra, they showed that a free locally compact Hausdorff étale groupoid with finite dynamic asymptotic dimension is amenable [[GWY17], Corollary 8.25]. Unlike their proof method, we adopt a direct method analogous to [[GWY22], Theorem A.9]; hence, we strengthen their result by removing the freeness assumption on groupoids.
Theorem 6.3.
Let $\mathcal{G}$ be a locally compact, Hausdorff, étale groupoid with finite dynamic asymptotic dimension. Then $\mathcal{G}$ is amenable.
Proof.
Assume that $\mathcal{G}$ has dynamic asymptotic dimension $d$. For any compact subset $K$ of $\mathcal{G}$ and any $\varepsilon>0$. Using local compactness, we expand $K$ slightly and write it as $K^{\prime}$, which is open and precompact. Replacing $K^{\prime}$ with $K^{\prime}\cup K^{\prime-1}\cup r(K^{\prime})\cup s(K^{\prime})$, we may assume that $K^{\prime}=K^{\prime}\cup K^{\prime-1}\cup r(K^{\prime})\cup s(K^{\prime})$, which is also open and precompact. Let $\{U_{0},\cdots,U_{d}\}$ be an open cover of $r(K^{\prime})\cup s(K^{\prime})$ with the properties in Proposition 5.8 for the open precompact set $K^{\prime}$ and error estimate $\varepsilon/d+2$. Replacing each $U_{i}$ with its intersection with $s(K^{\prime})\cup r(K^{\prime})$, we may assume that each $U_{i}$ is contained in $s(K^{\prime})\cup r(K^{\prime})\subseteq K^{\prime}$. Claim: $K^{\prime}=\bigcup\limits_{i}K_{i}$, where $K_{i}=\{k\in K^{\prime}\mid s(k),r(k)\in U_{i}\}$, $K_{i}$ is open, symmetric, precompact and has compact closure in $\mathcal{G}_{i}$. By Lemma 5.5 (ii) (iii), $s(K^{\prime}\cap r^{-1}(U_{i}^{(0)}))\subseteq U_{i}^{(1)}$ and $U_{i}^{(0)}\subseteq U_{i}^{(1)}\subseteq U_{i}$, thus $K^{\prime}\cap r^{-1}(U_{i}^{(0)})\subseteq K_{i}$. By Lemma 5.5 (i), $\{U_{0}^{(0)},\cdots,U_{d}^{(0)}\}$ covers $\mathcal{G}^{(0)}$. Therefore,
$\bigcup\limits_{i}K_{i}\supseteq\bigcup\limits_{i}K^{\prime}\cap r^{-1}(U_{i}^{(0)})$ $=K^{\prime}\cap r^{-1}(\bigcup\limits_{i}U_{i}^{(0)})$
$=K^{\prime}\cap r^{-1}(\mathcal{G}^{(0)})\supseteq K^{\prime}.$
Clearly, $K_{i}$ is open, precompact and has compact closure in $\mathcal{G}_{i}$, it suffices to show $K_{i}=K_{i}^{-1}$. Indeed, for any $k\in K_{i}$, namely, $k\in K^{\prime}$, $s(g),r(g)\in U_{i}$. Thus $k^{-1}\in K^{\prime}$ as $K^{\prime}=K^{\prime-1}$, and $r(k^{-1})=s(k),s(k^{-1})=r(k)\in U_{i}$, which implies that $k^{-1}\in K_{i}$.
Let $\mathcal{G}_{i}$ be an open subgroupoid of $\mathcal{G}$ generated by the set $K_{i}$, then $\mathcal{G}_{i}$ is amenable follows from Lemma 6.2. Therefore, for each $i$, let $\mu_{i}:\mathcal{G}_{i}\rightarrow[0,1]$ be a continuous, compactly supported function as in the definition of amenability with respect to the compact subset $\overline{K_{i}}$ and error estimate $\frac{\varepsilon}{d+2}$. Extending $\mu_{i}$ to the whole $\mathcal{G}$ by defining $\mu_{i}(x)=0$ when $x\notin\mathcal{G}_{i}$. We define $\mu:\mathcal{G}\rightarrow[0,1]$, $g\mapsto\sum\limits_{i=0}^{d}\phi_{i}(s(g))\mu_{i}(g)$. Clearly, $\mu$ is a well-defined continuous, compactly supported function in $\mathcal{G}$.
For any $x\in\mathcal{G}^{(0)}$,
$\sum\limits_{g\in\mathcal{G}_{x}}\mu(g)$ $=\sum\limits_{g\in\mathcal{G}_{x}}\sum\limits_{i=0}^{d}\phi_{i}(s(g))\mu_{i}(g)$
$=\sum\limits_{i=0}^{d}\phi_{i}(x)\sum\limits_{g\in(\mathcal{G}_{i})_{x}}\mu_{i}(g)$
$\leq\sum_{i=0}^{d}\phi_{i}(x)\leq 1,$
where the third inequality comes from the amenability of $\mathcal{G}_{i}$ and the fact that if $x\in U_{i}$, then $x\in K^{\prime}$ and $s(x)=r(x)=x\in U_{i}$, thus $x\in K_{i}\subseteq\mathcal{G}_{i}$, and if $x\notin U_{i}$, then $\phi_{i}(x)=0$ as supp ($\textnormal{supp }(\phi_{i})\subseteq U_{i}$. The last equality is due to Proposition 5.8.
For any $k\in K\subseteq K^{\prime}$,
$\left|1-\sum\limits_{g\in\mathcal{G}_{r(k)}}\mu(g)\right|$ $=\left|1-\sum\limits_{g\in\mathcal{G}_{r(k)}}\sum\limits_{i=0}^{d}\phi_{i}(s(g))\mu_{i}(g)\right|$
$=\left|\sum\limits_{i=0}^{d}\phi_{i}(r(k))(1-\sum\limits_{g\in(\mathcal{G}_{i})_{r(k)}}\mu_{i}(g))\right|$
$\leq\sum\limits_{i=0}^{d}\phi_{i}(r(k))\left|1-\sum\limits_{g\in(\mathcal{G}_{i})_{r(k)}}\mu_{i}(g)\right|$
$<\sum\limits_{i=0}^{d}\phi_{i}(r(k))\varepsilon=\varepsilon,$
where the fourth inequality follows from the amenability of $\mathcal{G}_{i}$ and the fact that if $r(k)\in U_{i}$, then $r(k)\in K_{i}$ and $r(r(k))=r(k)$, if $r(k)\notin U_{i}$, then $\phi_{i}(r(k))=0$. The fifth equality is due to the partition of unity.
For any $k\in K\subseteq K^{\prime}$,
$\sum\limits_{g\in\mathcal{G}_{r(k)}}\left|\mu(g)-\mu(gk)\right|$ $=\sum\limits_{g\in\mathcal{G}_{r(k)}}\left|\sum\limits_{i=0}^{d}\left(\phi_{i}(s(g))\mu_{i}(g)-\phi_{i}(s(gk))\mu_{i}(gk)\right)\right|$
$\leq\sum\limits_{i=0}^{d}\sum\limits_{g\in\mathcal{G}_{r(k)}}\phi_{i}(r(k))\left|\mu_{i}(g))-\mu_{i}(gk)\right|+\mu_{i}(gk)\left|\phi_{i}(r(k))-\phi_{i}(s(k))\right|$
$=\sum\limits_{i=0}^{d}\phi_{i}(r(k))\sum\limits_{g\in(\mathcal{G}_{i})_{r(k)}}\left|\mu_{i}(g)-\mu_{i}(gk)\right|$
$+\sum\limits_{i=0}^{d}\left|\phi_{i}(r(k))-\phi_{i}(s(k))\right|\sum\limits_{g\in(\mathcal{G}_{i})_{r(k)}}\mu_{i}(gk)$
$<\sum\limits_{i=0}^{d}\phi_{i}(r(k))\cdot\frac{\varepsilon}{d+2}+(d+1)\cdot\frac{\varepsilon}{d+2}\cdot 1=\varepsilon,$
where the fourth inequality comes from the amenability of $\mathcal{G}_{i}$ and Proposition 5.8. ∎
The following result is inspired by [[Oza12], Theorem 1], which demonstrated that subexponential asymptotic dimension growth for a metric space implies property A. Clearly, slow asymptotic dimension growth also implies property A as it is a subexponential growth. Since asymptotic dimension growth for a discrete metric space of bounded geometry is equivalent to the dynamic asymptotic dimension growth for the corresponding coarse groupoid (Theorem 4.12), and as property A for a discrete metric space of bounded geometry is equivalent to amenability of the corresponding coarse groupoid [[STY02], Theorem 5.3], we obtain a generalization of the result of Corollary 4.15 in the case of slow growth.
Theorem 6.4.
Let $\mathcal{G}$ be a $\sigma$-compact étale groupoid with compact unit space, and with dynamic asymptotic dimension growth $f\preceq x^{\alpha}(0<\alpha<1)$. Then $\mathcal{G}$ is amenable.
Proof.
For any compact subset $K$ of $\mathcal{G}$ and any $\varepsilon>0$, there exists $R>0$ such that $K\subset K^{\prime}=\{z\in\mathcal{G}\mid\ell(z)<R\}$ as the length function $\ell$ is controlled. Note that $K^{\prime}$ is open and precompact, $K^{\prime}=K^{\prime-1}$, $s(K^{\prime})\cup r(K^{\prime})=\mathcal{G}^{(0)}\subseteq K^{\prime}$. By Proposition 5.1, there exists a constant $c$ depending on parameters $\alpha,R,\varepsilon$, a positive integer $p$, and an open cover $\{U_{0},\cdots,U_{f([(R+1)^{c}]R+1)}\}$ of $\mathcal{G}^{(0)}$ with the three properties. Let $\mathcal{G}_{i}$ be an open precompact subgroupoid generated by the set $K_{i}=\{k\in K^{\prime}\mid s(k),r(k)\in U_{i}\}$ and error estimate $\frac{\varepsilon}{2}$. Then by Lemma 6.2, $\mathcal{G}_{i}$ is amenable.
For each $i$, let $\mu_{i}:\mathcal{G}_{i}\rightarrow[0,1]$ be a continuous, compactly supported function as in the definition of amenability with respect to the compact subset $\overline{K_{i}}$ and error estimate $\frac{\varepsilon}{2}$. Extending $\mu_{i}$ to the whole $\mathcal{G}$ by defining $\mu_{i}(x)=0$, if $x\notin\mathcal{G}_{i}$. We define $\mu:\mathcal{G}\rightarrow[0,1],g\mapsto\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(s(g))\mu_{i}(g)$, which is well-defined continuous, compactly supported function on $\mathcal{G}$.
For any $x\in\mathcal{G}^{(0)}$,
$\sum\limits_{g\in\mathcal{G}_{x}}\mu(g)$ $=\sum\limits_{g\in\mathcal{G}_{x}}\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(s(g))\mu_{i}(g)$
$=\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(x)\sum\limits_{g\in(\mathcal{G}_{i})_{x}}\mu_{i}(g)$
$\leq\sum_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(x)=1,$
where the third inequality comes from the amenability of $\mathcal{G}_{i}$, and the fourth one is due to the partition of unity.
For any $k\in K\subseteq K^{\prime}$,
$\left|1-\sum\limits_{g\in\mathcal{G}_{r(k)}}\mu(g)\right|$ $=\left|1-\sum\limits_{g\in\mathcal{G}_{r(k)}}\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(s(g))\mu_{i}(g)\right|$
$=\left|\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(r(k))(1-\sum\limits_{g\in(\mathcal{G}_{i})_{r(k)}}\mu_{i}(g))\right|$
$\leq\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(r(k))\left|1-\sum\limits_{g\in(\mathcal{G}_{i})_{r(k)}}\mu_{i}(g)\right|$
$<\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(r(k))\cdot\frac{\varepsilon}{2}=\frac{\varepsilon}{2}<\varepsilon,$
where the fourth inequality comes from the amenability of $\mathcal{G}_{i}$, and the fifth equality is due to the partition of unity.
For any $k\in K\subseteq K^{\prime}$,
$\sum\limits_{g\in\mathcal{G}_{r(k)}}\left|\mu(g)-\mu(gk)\right|$ $\leq\sum\limits_{g\in\mathcal{G}_{r(k)}}\left|\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\left(\phi_{i}^{p}(s(g))\mu_{i}(g)-\phi_{i}^{p}(s(gk))\mu_{i}(gk)\right)\right|$
$\leq\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\sum\limits_{g\in\mathcal{G}_{r(k)}}\phi_{i}^{p}(r(k))\left|\mu_{i}(g)-\mu_{i}(gk)\right|+\mu_{i}(gk)\left|\phi_{i}^{p}(r(k))-\phi_{i}^{p}(s(k))\right|$
$=\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(r(k))\cdot\sum\limits_{g\in(\mathcal{G}_{i})_{r(k)}}\left|\mu_{i}(g)-\mu_{i}(gk)\right|$
$+\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\left|\phi_{i}(r(k))-\phi_{i}(s(k))\right|\cdot\left|\sum\limits_{j=0}^{p-1}\phi_{i}(r(k))^{j}\phi_{i}(s(k))^{p-1-j}\right|\cdot\sum\limits_{g\in(\mathcal{G}_{i})_{r(k)}}\mu_{i}(gk)$
$<\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\phi_{i}^{p}(r(k))\cdot\frac{\varepsilon}{2}+p\cdot\sum\limits_{i=0}^{f([(R+1)^{c}]R+1)}\left|\phi_{i}(r(k))-\phi_{i}(s(k))\right|$
$<\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon,$
where the third equality comes from the equality $\left|a^{p}-b^{p}\right|=\left|a-b\right|\cdot\left|\sum\limits_{j=0}^{p-1}a^{j}b^{p-1-j}\right|$ for all $a,b\geq 0$, the fourth inequality is due to the amenability of $\mathcal{G}_{i}$, and the final inequality follows from Proposition 5.1(3). ∎
Example 6.5.
Let $f:\mathbb{R}^{+}\rightarrow\mathbb{N}$ be a strictly increasing function such that $f\preceq x^{\alpha}(0<\alpha<1)$, and let $X=\bigoplus\limits_{n=1}^{\infty}f(n)\mathbb{Z}$, endowed with the direct sum metric. Then $X$ is a discrete metric space with asymptotic dimension growth $f$, a subexponential growth. Note that $X$ has bounded geometry since the sequence $\{f(n)\}$ is (strictly) increasing and every element of $X$ has only finitely many nonzero terms, thus $X$ has property $A$. Hence, the corresponding coarse groupoid $G(X)$ is amenable by Corollary 4.15.
Acknowledgments
We thank Jiawen Zhang for many valuable suggestions on asymptotic dimension growth and Jianchao Wu for useful discussions on coarse groupoids.